### v1

# M3-20M: A Large-Scale Multi-Modal Molecule Dataset for AI-driven Drug Design and Discovery

Abstract
This paper introduces M3-20M, a large-scale Multi-Modal Molecular dataset that contains over 20 million molecules. Designed to support AI-driven drug design and discovery, M3-20M is 71 times more in the number of molecules than the largest existing dataset, providing an unprecedented scale that can highly benefit training or fine-tuning large (language) models with superior performance for drug design and discovery. This dataset integrates one-dimensional SMILES, two-dimensional molecular graphs, three-dimensional molecular structures, physicochemical properties, and textual descriptions collected through web crawling and generated by using GPT-3.5, offering a comprehensive view of each molecule. To demonstrate the power of M3-20M in drug design and discovery, we conduct extensive experiments on two key tasks: molecule generation and molecular property prediction, using large language models including GLM4, GPT-3.5, and GPT-4. Our experimental results show that M3-20M can significantly boost model performance in both tasks. Specifically, it enables the models to generate more diverse and valid molecular structures and achieve higher property prediction accuracy than the existing single-modal datasets, which validates the value and potential of M3-20M in supporting AI-driven drug design and discovery. The dataset is available at https://github.com/bz99bz/M-3.
keywords:
American Chemical Society,
[…]
{tocentry}
1 Introduction
Drug design and discovery is an important field in pharmacy, which aims to identify new therapeutic compounds and optimize their properties for clinical use 1, 2, 3, 4, 5. This process involves the prediction of molecule-target interactions 6, 7, 8, 9 and molecular properties 10, 11, 12, 13, and the design or generation of targeted molecules 14, 15, 16, 17, 18, which are fundamental for developing effective and safe drugs. In recent years, various models trained or pre-trained and fine-tuned using massive data, such as 19, 20, 21, 22, 23, 24, 25, 26, 27, have emerged as powerful tools in this domain. These models leverage large datasets and sophisticated algorithms to generate new compounds, predict molecular structures, and analyze their biochemical properties, significantly advancing the process of drug design and discovery. However, most existing molecular datasets for model training are limited to single-modality, failing to capture the characteristics of molecules completely, thus hampering the building of powerful models for drug design and discovery.
Recent works 23, 28, 29 have been paying increasing attention to multi-modal molecule datasets. Zeng 29 introduced a deep-learning system that bridges molecular structures and biomedical texts, achieving comprehension comparable to human professionals. Liu et al. 28 developed a multi-modal model for text-based retrieval and editing of molecular structures, facilitating advanced molecular text analysis. Liu et al. 23 presented a multi-modal large language model for molecular science that integrates graphs, images, and texts to support various downstream tasks. However, datasets used in these works still suffer from significant limitations. On the one hand, these datasets are typically small in size, encompassing a limited chemical space that restricts the generalization power of the trained or tuned models. On the other hand, the absence of complete modalities of molecular data undermines the performance of computational models.
To address these limitations, this paper presents a new and large-scale multi-modal integrated molecule dataset called M3-20M for AI-driven drug design and discovery. M3-20M contains more than 20 million molecules with their SMILES strings, 2D graphs, 3D structures, physicochemical properties, and textual descriptions, providing unprecedented data scale, diversity, and comprehensiveness for training models of superior performance for various drug design and discovery downstream tasks. Concretely, M3-20M is 71 times more in the number of molecules than the largest existing dataset28. It collects physicochemical properties from the PubChem database 30 and enriches the description texts of molecules by both web crawling and GPT-3.5 generation. Therefore, each molecule’s SMILES corresponds to its 2D graph and 3D structure, physicochemical properties, and description texts. Figure 1 shows some examples from our M3-20M dataset.
In summary, M3-20M stands out of the existing datasets in at least the following three aspects: 1) Large scale: it contains over 20 million molecules, which is the largest open-access multi-modal molecule dataset for AI-driven drug design and discovery, to the best of our knowledge. 2) Comprehensive modalities: M3-20M boasts a more complete range of modalities, including one-dimensional molecular SMILES strings, two-dimensional molecular graphs, three-dimensional molecular structures, physicochemical properties, and textual descriptions. Such multi-modal data offers a holistic perspective of each molecule, benefiting model training and tuning for drug design and discovery. Additionally, we offer tools to generate 2D molecular graph images that facilitate compound patent retrieval, and to crawl PubMed for research articles related to specific molecules, thereby enriching textual descriptions. 3) Supporting various tasks: M3-20M can support model training and tuning for various downstream drug design and discovery tasks, including molecule generation, molecular property prediction, lead optimization, and virtual screening. It also can aid in tasks such as pharmacokinetics modeling and drug-target interaction prediction. Moreover, we construct seven multimodal sub-datasets, QM9-MM, MOSES-MM, BACE-MM, BBBP-MM, HIV-MM, ClinTox-MM, and Tox21-MM, to facilitate more accurate and robust molecular property prediction.
To demonstrate the power of M3-20M in supporting drug design and discovery, extensive experiments on two basic tasks, molecule generation and molecular property prediction, are conducted using large language models including GLM4 31, GPT-3.5 32, and GPT-4 33. The experimental results show that M3-20M can significantly boost model performance in both tasks. Concretely, with M3-20M, the above models can generate more diverse and valid molecular structures, and achieve higher property prediction accuracy than with existing single-modal datasets, thus verifying the value and potential of M3-20M in supporting AI-driven drug design and discovery.
Figure 1: Some examples from our M3-20M dataset.
2 Related Work
Multi-modal molecular data is becoming increasingly important in artificial intelligence and molecular biology. which can provide more comprehensive and rich information, bringing new opportunities and challenges to drug discovery, molecular property prediction, and generation tasks. For example, Wang et al. 34 pointed out that the use of multi-modal data, including sequence, structure, text etc., can improve the accuracy of drug discovery. However, the challenges faced by multi-modal data cannot be ignored, as different data types often necessitate distinct methods for organization, storage and processing as well as learning. Additionally, ensuring data consistency and integrity also poses a critical issue, with various forms of data potentially exhibiting semantic gaps or inconsistencies. Therefore, there is an urgent need to overcome the above issues to create a large-scale multi-modal integrated dataset for model training or tuning in drug design and discovery. In what follows, we briefly review the related work in three perspectives, molecular representation, single and two-modality molecule datasets, and multi-modality molecule datasets.
2.1 Molecular representations
Molecules can be represented by various modalities of information 35, 36, including one-dimensional sequences, two-dimensional graphs or images, three-dimensional structures, and molecular properties as well as textual descriptions. One-dimensional sequences primarily consist of SMILES strings, molecular fingerprints, and algebraic topology-based characterizations. Two-dimensional representations include molecular graphs and images generated using toolkits such as RDKit 37 or Open Babel 38. Three-dimensional structures are categorized into molecular graphs and grids 39, containing atomic coordinates that provide detailed spatial information. Molecular textual descriptions offer additional information derived from experiments. Single-modal data is a reduction of the complete structural information of the molecules, which means the representations learned from single-modal suffer from information loss. For example, two SMILES may represent the same molecule or a two-dimensional molecular graph cannot distinguish between the conformations. Therefore, integrating multi-modal information is crucial for molecular modeling and model training in drug design and discovery.
2.2 Single-modal and two-modal molecule datasets
Most existing molecule datasets are limited to single or dual modalities, restricting the generalization power of trained models 40, 41, 42, 43, 44. These datasets typically have only one-dimensional (1D) SMILES strings, or two-dimensional (2D) structures, or three-dimensional (3D) structures, and usually lack textual molecular descriptions. For instance, MOSES 45 comprises 4,591,276 molecules filtered by molecular weight, ranging from 250 to 350 Daltons, and contains solely one-dimensional SMILES representations. Similarly, ZINC 46, which serves as a repository of commercially available compounds for virtual screening, contains over 1.3 billion molecules obtained from 310 catalogs across 150 vendors, encompassing 2D graphs and, for the majority, 3D structures. Zeng et al. 29 presented a deep learning system that connects molecular structures with biomedical texts, demonstrating a level of comprehension comparable to that of human experts.
2.3 Multi-modal molecule datasets
With the development of AI-driven drug design and discovery, multi-modal data have become increasingly important in recent years. Up to now, only a few datasets have been introduced to address this need, each with distinct purposes and contributions.
Liu et al. 28 constructed PubChemSTM, a dataset containing over 280,000 chemical structure–text pairs. Despite PubChemSTM’s ability to generalize novel biochemical concepts across various benchmarks, its dataset volume remains insufficient. Subsequently, the igcdata dataset 23 was proposed, which comprises 220,000 descriptions sourced from PubChem and ChEBI-20 47. Although these two datasets provide valuable resources, their relatively small scale restricts their utility for training high-performance models. Beyond scale, these two datasets also lack the diversity and completeness of molecular modalities necessary for robust model development. Issues such as sparse or overly brief textual descriptions further exacerbate this limitation, leading to suboptimal training outcomes of drug design and discovery models.
To address the limitations of existing datasets and support the training or tuning of high-performance models for drug design and discovery, we present the M3-20M dataset, which significantly expands the scale and scope of multi-modal molecular data. Table 1 offers a qualitative comparison of our dataset against existing multi-modal datasets across seven dimensions, including the number of molecules contained, whether consisting of 3D structures, 2D graphs, PubChem CID, physicochemical properties and downstream task datasets, and the number of description texts. From Table 1, we can see that our dataset M3-20M has the largest number of molcuels and covers the most comprehensive molecular data modalities.
Property QM948 PCdes29 PubChemSTM28 igcdata23 M3-20M (Ours)
SMILES 134K 1.5K 280K 220K 20M
3D structures $\checkmark$ × × × $\checkmark$
2D graphs $\times$ $\times$ $\checkmark$ × $\checkmark$
2D images $\times$ $\times$ × $\checkmark$ $\checkmark$
PubChem CID $\times$ × $\checkmark$ × $\checkmark$
Physicochemical properties $\times$ × × × $\checkmark$
Downstream task datasets $\times$ × × × $\checkmark$
#Text descriptions $\times$ 1.5K 280K 220K 20M
Table 1: Comparison between our dataset M3-20M and existing multi-modal datasets.
3 Dataset Construction
This section introduces the dataset construction process and some statistic information of the dataset. Molecular data is collected from existing databases PubChem, ZINC and QM9, and then processed to get multiple representations of molecules, including SMILES strings, 2D and 3D structures, physicochemical properties, and textual descriptions. Notably, we enhance molecular textual descriptions by generating using GPT-3.5 with an expert scoring mechanism to ensure that these descriptions are both scientifically accurate and practically useful. All the data are integrated into a comprehensive dataset with over 20 million molecules.
Figure 2: The construction pipeline of our multi-modal molecular dataset M3-20M.
3.1 Data Collection
Figure 2 shows the pipeline for constructing our multi-modal molecular dataset M3-20M. To obtain molecules and the corresponding molecular graphs and 3D structures, we first collect molecules from the PubChem, ZINC, and QM9 databases for their extensive molecular coverage and reliable data quality. Then, we utilize the Chem function of RDKit to extract atomic features and chemical bond characteristics and convert them into 2D molecular graphs. Additionally, we employ the PubChem API to batch download SDF files containing the 3D structures of molecules and use the GetAtomPosition function of RDKit to calculate the 3D coordinates of the central atoms.
Textual descriptions serve as a bridge for large language models to comprehend multi-modal molecules. The remarkable recognition accuracy and robustness of the CLIP 49 model exemplify the benefits of effectively linking textual descriptions with other modality information of molecules. Similarly, since textual descriptions of molecules facilitate a deeper understanding of chemical information, we meticulously construct these descriptions in the dataset using three approaches: 1) For the molecules that have textual descriptions in PubChem, we directly extract the texts from PubChem. 2) For the molecules that do not have textual descriptions but have physicochemical properties in PubChem, we transform these physicochemical property values into textual descriptions, which will be detailed later. 3) For those without physicochemical property values or not present in the PubChem database, we generate their textual descriptions by GPT-3.5. We present the details in the following sections.
Initially, we utilized textual data from PubChem; however, only a subset of molecules in PubChem contains associated textual descriptions. The sparsity of text information leads to poor performance of the trained or tuned models. So we resort to text augmentation. To generate textual descriptions for molecules having physicochemical property values in PubChem, we extract 26 key physicochemical properties and their values from PubChem, including Molecular Weight, XLogP3, Hydrogen Bond Donor Count, Hydrogen Bond Acceptor Count, Rotatable Bond Count, Exact Mass, Monoisotopic Mass, Topological Polar Surface Area, Heavy Atom Count, Formal Charge, Complexity, Isotope Atom Count, Defined Atom Stereocenter Count, Undefined Atom Stereocenter Count, Defined Bond Stereocenter Count, Undefined Bond Stereocenter Count, Covalently-Bonded Unit Count, Canonicalization Status, Physical Description, Color/Form, Odor, Boiling Point, Melting Point, Flash Point, and Solubility Density.
Concretely, we search for these properties using the compound identifier (CID) in PubChem, ensuring all data complied with licensing requirements and utilizing the open interface provided by PubChem. After collecting 19,175,245 pieces of raw data, we transform the properties of each molecule into structured descriptions using the template “property name is a specific value”. Here, “property name” indicates any of the 26 properties, and “specific value” means a concrete value of the property.
Furthermore, we expand seven downstream task datasets with multi-modal information from our dataset by associating the PubChem CIDs of molecules, including molecule generation datasets MOSES and QM9, and molecular property prediction datasets such as BBBP, BACE, HIV, Tox21, and ClinTox. This expansion enables model enhancement with multi-modal information for both molecule generation and property prediction tasks. The expanded datasets are denoted as MOSES-MM, QM9-MM, BBBP-MM, BACE-MM, HIV-MM, Tox21-MM, and ClinTox-MM, respectively.
3.2 Text Description Generation by GPT-3.5
As mentioned above, for molecules lacking physicochemical properties or not present in PubChem, we generate molecular textual descriptions as a molecule captioning task by using the GPT-3.5 API. Figure 3 illustrates the process of molecular description text generation. Initially, we use the following prompt to guide GPT-3.5 in generating descriptions based solely on SMILES strings: “You are an expert chemist. Given the molecular SMILES, your task is to provide a detailed description of the molecule using your experienced chemical knowledge.”
Figure 3: Molecular text description generation by GPT-3.5.
However, this approach results in issues such as incorrect naming of SMILES strings and misidentifying functional groups. To address these problems, inspired by the CO-STAR (Context-Objective-Style-Tone-Audience-Response) framework, we rewrite the original prompt as follows to enhance the output quality: “You are an expert chemist with a comprehensive understanding of molecular structures and chemistry. Using your extensive knowledge, please describe the molecule represented by the following SMILES string. Focus on its structural features, potential chemical properties, and any known applications or reactions it may participate in. Use only the information inferred from the SMILES representation; do not create or infer any hypothetical or novel names for the molecule.”
Utilizing this enhanced prompt, we generated 1,073,845 descriptions using GPT-3.5, which is only 0.934% of the total descriptions. The proportion of synthetic data to real data is very low and will not have a negative impact on real data. The following section elaborates on the quality control measures implemented through the Human Expert Scoring Mechanism to ensure the precision and reliability of the generated descriptions.
3.3 Generated Text Quality Control by Human Expert Scoring
To ensure the reliability and utility of the generated textual descriptions, we implement an expert scoring mechanism inspired by some recent works 50, 51. We recruit six undergraduate and postgraduate chemistry students from top China universities, all having excellent academic records. They evaluate the accuracy and usefulness of the generated texts, ensuring that the textual descriptions are both scientifically valid and practically useful.
This mechanism assesses the textual descriptions from four dimensions: accuracy, effectiveness, comprehensiveness, and simplicity, as established in previous works on natural language generation and evaluation 52, 53. Each description can earn a maximum score of 10 points. The scoring system is structured as follows:
  1. 1.
    Accuracy (Maximum 5 points)
    1. (a)
      Correct Description (5 points): The molecular description is entirely accurate, containing no scientific errors. This includes correct usage of chemical terms, accurate structural representation, and proper property identification.
    2. (b)
      Slight Errors (2-4 points): Minor inaccuracies are present, such as small structural mistakes or minor property deviations. These errors should not impede overall comprehension. Notably, any incorrect identification of functional groups results in a 1 point deduction.
    3. (c)
      Serious Errors (0 points): Major inaccuracies are present, such as entirely incorrect chemical terminology or wrong molecular names, stripping the description of scientific value.
  2. 2.
    Effectiveness (Maximum 2 points)
    1. (a)
      Highly Valid (2 points): The description highlights key molecular properties, such as reactivity or pharmacodynamics, and may include useful applications or future prospects.
    2. (b)
      Moderately Effective (1 point): While the description includes valid information, it may lack detailed analysis or specific application scenarios.
    3. (c)
      Ineffective Description (0 points): The description fails to convey crucial properties or applications and lacks practical utility.
  3. 3.
    Comprehensiveness (Maximum 2 points)
    1. (a)
      Very Comprehensive (2 points): The description extensively covers multiple aspects of the molecule, including structure, physicochemical properties, reactivity, and potential applications.
    2. (b)
      Fairly Comprehensive (1 point): Several aspects of the molecule are described, though some areas may lack depth.
    3. (c)
      Incomplete (0 points): The description is limited to one aspect of the molecule and significantly lacks breadth.
  4. 4.
    Simplicity (Maximum 1 point)
    1. (a)
      Concise (1 point): Succinct, clearly articulated, and free of redundant information.
    2. (b)
      Not Concise (0 points): Verbose and repetitive, detracting from clarity and conciseness.
Each text description is evaluated based on a maximum score of 10 points. For every 100 generated molecules, 10% are randomly selected for expert scoring. Descriptions scoring above 5 points are considered qualified, while those below 5 points are regenerated. Descriptions with significant inaccuracies (0 points in accuracy) will also be regenerated.
4 Statistic Results
This section presents some statistic information of the M3-20M dataset to highlight its comprehensive annotations, diverse textual descriptions and physicochemical properties, which can significantly enhance its utility for molecular research and AI-driven drug discovery. By comparing it with existing datasets such as PubChem, we emphasize the enhancements in both scale and quality that M3-20M offers. The section further delves into word distribution — a key feature of textual data, and explores correlations among physicochemical properties through advanced statistical techniques like Principal Component Analysis (PCA). Additionally, we outline the robust maintenance protocols established to ensure our dataset’s accuracy and usability for the research community. These findings demonstrate the dataset’s value as a fundamental resource for molecular representation learning and related applications.
4.1 Textual Descriptions
Figure 4 shows a comparison of molecule descriptions between PubChem and our M3-20M. The PubChem description bar is relatively small, reaching only 360,133, while the M3-20M annotation description bar towers over it, marking a count of 20,249,090. This enhancement is critical for AI-driven drug design and discovery, as the increased volume and details of descriptions facilitate more comprehensive analysis and model training.
Figure 4: Comparison of molecule descriptions: PubChem vs. our M3-20M
4.2 Word Distribution in Description Texts
Figure 5 presents the top 20 most frequently occurring words in the textual descriptions of our dataset, plotted alongside their respective frequencies to emphasize their relative prominence. The most commonly used terms, such as ’natural,’ ’product,’ and ’acid,’ appear tens of thousands of times, underscoring their central role in describing molecular properties and functions. Additional notable terms, including ’related,’ ’organisms,’ and ’amino,’ further highlight the biological and chemical contexts associated with these molecules. This visualization of word frequencies provides valuable insights into molecular datasets’ linguistic patterns and descriptive focus.
Figure 5: Top 20 most frequent words in textual descriptions.
4.3 Physicochemical Properties
Table 2 and Figure 6 provide a detailed overview of the specific names, corresponding quantities, and reference of the 26 key physicochemical properties extracted from PubChem. Of the 26 key physicochemical properties, 18 are computed using professional software  [^1] PubChem 2.2; Cactvs 3.4.8.18, while the remaining 8 represent experimental values sourced from the CAMEO Chemicals database and Hazardous Substances Data Bank (HSDB). Notably, 18 of these properties exhibit quantities reaching tens of millions, underscoring our dataset’s extensive coverage and comprehensiveness.
Physicochemical Properties Number Reference Molecular Weight 19174000 Computed by PubChem 2.2 XLogP3 16340203 Computed by XLogP3 3.0 Hydrogen Bond Donor Count 19173687 Computed by Cactvs 3.4.8.18 Hydrogen Bond Acceptor Count 19173559 Computed by Cactvs 3.4.8.18 Rotatable Bond Count 19173464 Computed by Cactvs 3.4.8.18 Exact Mass 19173398 Computed by PubChem 2.2 Monoisotopic Mass 19173327 Computed by PubChem 2.2 Topological Polar Surface Area 19173287 Computed by Cactvs 3.4.8.18 Heavy Atom Count 19173259 Computed by PubChem Formal Charge 19173222 Computed by PubChem Complexity 19173210 Computed by Cactvs 3.4.8.18 Isotope Atom Count 19173193 Computed by PubChem Defined Atom Stereocenter Count 19173170 Computed by PubChem Undefined Atom Stereocenter Count 19173153 Computed by PubChem Defined Bond Stereocenter Count 19173147 Computed by PubChem Undefined Bond Stereocenter Count 19173137 Computed by PubChem Covalently-Bonded Unit Count 19173132 Computed by PubChem Compound Is Canonicalized 19173130 Computed by PubChem Physical Description 2031 CAMEO Chemicals database Color/Form 246 Hazardous Substances Data Bank (HSDB) Odor 100 Hazardous Substances Data Bank (HSDB) Boiling Point 130 CAMEO Chemicals database Melting Point 633 CAMEO Chemicals database Flash Point 35 CAMEO Chemicals database Solubility 10813 CAMEO Chemicals database Density 170 CAMEO Chemicals database
Table 2: The 26 key physicochemical properties and their corresponding numbers and references.
Figure 6: Physicochemical property distribution.
Figure 7: Correlation matrix of physicochemical properties.
Figure 8: Visualization of molecular space with principal component analysis (PCA).
As shown in Figure 7, we first calculate the correlation matrix of all physicochemical properties and draw a heat map. We find that Molecular Weight and XLogP3 are the most suitable because they exhibit strong correlations with several other key properties, indicating their ability to capture significant variance in the dataset. Thus, we select these two properties as the first and second principal components for Principal Components Analysis (PCA) and visualize the molecular space of randomly sampled 10000 molecules as shown in Figure 8.
4.4 Dataset Maintenance
To ensure the accessibility and accuracy of the M3-20M dataset, we have organized a dedicated maintenance team consisting of six members, including three doctoral students and three master’s students. This team will perform regular maintenance twice a week, correcting any errors and updating the dataset as needed. This rigorous maintenance protocol supports the community’s exploration of molecular representation and downstream molecular tasks and facilitates the training of large models with improved versatility and robustness.
4.5 Supporting Various Downstream Tasks
The M3-20M dataset supports model training and tuning for various downstream drug design and discovery tasks. While general large language models have achieved remarkable success in natural language processing, their performance in specialized fields like drug design and discovery still has significant room for improvement. Figure 9 illustrates how to use the M3-20M dataset to enhance large language models (LLMs) or train models for downstream drug design and discovery tasks, which consists of three paradigms: prompting, finetuning, and training from scratch.
Prompt engineering involves carefully constructing input prompts to guide a general large language model (LLM) in generating more accurate outputs tailored to drug design and discovery. By leveraging our M3-20M dataset, prompt engineering allows the model to quickly adapt to this field without altering its internal parameters. Fine-tuning refers to further training the LLMs on the M3-20M dataset, adjusting certain model parameters based on existing pre-training. This process enables the model to better align with the specific characteristics and needs of drug design and discovery while retaining its general language knowledge. Training from scratch entails building and training a language model on drug design and discovery domain data. This approach necessitates a large volume of training data, which our M3-20M dataset can adequately provide. By utilizing our M3-20M dataset through prompt engineering, fine-tuning, and training from scratch, the general LLM’s drug design and discovery capabilities can be enhanced from multiple perspectives and to varying degrees.
M3-20M also can aid in tasks such as pharmacokinetics modeling and drug-target interaction prediction. Moreover, we construct seven multimodal sub-datasets, QM9-MM, MOSES-MM, BACE-MM, BBBP-MM, HIV-MM, ClinTox-MM, and Tox21-MM, to facilitate more accurate and robust molecular property prediction. We also provide tools for generating molecular images to support compound patent searches and for crawling PubMed to find articles related to specific molecules, thereby enhancing textual descriptions.
Figure 9: Paradigms of model training and finetuning using our dataset.
5 Dataset Evaluation
In this section, we evaluate the effectiveness of the M3-20M dataset on two fundamental downstream tasks: molecule generation and molecular property prediction. We primarily perform prompt tuning and in-context learning on the closed-source large language models GLM4, GPT-3.5, and GPT-4, while also conducting fine-tuning on the open-source model Llama3-8b. Additionally, our dataset can be applied to other tasks, including name prediction, reaction prediction, retrosynthesis, text-based molecule design, and molecule captioning. Our findings demonstrate that M3-20M significantly enhances model performance across these molecular tasks.
5.1 Molecule Generation
We utilize the GLM-4, GPT-3.5, and GPT-4 APIs to evaluate the power of M3-20M in molecule generation. Through in-context learning (ICL) 54, we provide 10 randomly selected multi-modal examples (including SMILES strings of molecules, three-dimensional atomic coordinates, and descriptions) and single-modal examples (containing only SMILES), respectively. Subsequently, 100 molecules are generated randomly for each setup.
We analyze the performance of models trained on single-modal and multi-modal datasets in molecule generation using three metrics: validity, uniqueness, and novelty. The results, presented in Table 3, demonstrate the superiority of multi-modal datasets across all metrics and models.
Validity measures the proportion of chemically valid generated molecules. Multi-modal datasets consistently outperform single-modal ones. For instance, GLM4’s validity improves from 72.73% to 85.37%, GPT-3.5’s from 76.01% to 84.8%, and GPT-4’s from 92.3% to 97.99%.
Uniqueness assesses the diversity of generated molecules. Multi-modal datasets show significant improvements: GLM4’s uniqueness rises from 68.17% to 82.25%, GPT-3.5’s from 46.46% to 93.86%, and GPT-4’s from 58.64% to 70.17%.
Novelty evaluates the proportion of generated novel molecules. Multi-modal datasets show substantial improvements, with GLM4’s novelty increasing from 97.90% to 98.10%, GPT-3.5’s from 85.72% to 96.51%, and GPT-4’s from 90.66% to 98.95%.
Validity(%)↑ Uniqueness(%)↑ Novelty(%)↑
Single-modal Multi-modal Single-modal Multi-modal Single-modal Multi-modal
GLM4 72.73($\pm$1.73) 85.37($\pm$1.58) 68.17($\pm$2.94) 82.25($\pm$2.95) 97.90($\pm$0.98) 98.10($\pm$1.23)
GPT-3.5 76.01($\pm$5.63) 84.8($\pm$0.6) 46.46($\pm$3.23) 93.86($\pm$1.49) 85.72($\pm$0.67) 96.51($\pm$1.57)
GPT-4 92.3($\pm$3.85) 97.99($\pm$0.41) 58.64($\pm$0.36) 70.17($\pm$4.92) 90.66($\pm$1.14) 98.95($\pm$0.34)
Table 3: Molecule generation performance comparison: multi-modal data vs. single-modal data.
Furthermore, Table 4 summarizes the molecular generation performance of our M3-20M dataset using MOSES-MM as the benchmark. Multi-modal data enhances SNN/Test performance for GPT-3.5 and GPT-4, showing increased similarity to the test set. For Frag/Test, multi-modal data boosts GLM4 to 0.27% and GPT-4 to 0.20%. In Scaf/Test, all models improve with multi-modal data; GLM4 rises to 0.07%, while GPT-3.5 and GPT-4 each improve to 0.01%. Regarding Filters, GLM4’s pass rate increases to 0.80% with multi-modal data, while GPT-3.5 maintains 1.0%.
Metrics like logP, SA, QED, and molecular weight are assessed using the Wasserstein-1 distance to the MOSES-MM test set, where lower values indicate better performance. Multi-modal data significantly reduces logP distance, especially for GPT-3.5 (0.99) and GPT-4 (0.92), indicating superior predictions. In SA, multi-modal data improves scores for GPT-3.5 and GPT-4, reducing distances to 0.57 and 0.54, respectively. QED scores improve with multi-modal data, reaching 0.12 for both models. For molecular weight, predictions improve for GLM4, reducing the distance to 105.72.
These findings demonstrate that our multi-modal dataset significantly enhances molecule generation models across nearly all metrics, highlighting the benefit of integrating diverse molecular representations. The results support the widespread use of multi-modal data in molecular research to develop more robust and versatile generative models.
GLM4 GPT-3.5 GPT-4 SNN/Test(%)↑ Single-modal 0.49(±0.03) 0.53(±0.09) 0.49(±0.07) Multi-modal 0.44(±0.09) 0.56(±0.04) 0.52(±0.01) Frag/Test(%)↑ Single-modal 0.24(±0.02) 0.44(±0.12) 0.19(±0.08) Multi-modal 0.27(±0.14) 0.32(±0.01) 0.20(±0.01) Scaf/Test(%)↑ Single-modal 0.04(±0.05) 6.90e-4(±6.90e-4) 2.01e-3(±4.13e-4) Multi-modal 0.07(±0.09) 0.01(±0.01) 0.01(±4.99e-3) Filters(%)↑ Single-modal 0.73(±0.09) 0.8(±0.0) 1.0(±0.0) Multi-modal 0.80(±1.11e-16) 1.0(±0.0) 0.9(±0.10) logP(↓) Single-modal 1.35(±0.17) 1.63(±0.32) 1.27(±0.31) Multi-modal 1.44(±0.28) 0.99(±0.05) 0.92(±0.11) SA(↓) Single-modal 0.83(±0.22) 0.63(±0.11) 0.60(±0.14) Multi-modal 1.00(±0.10) 0.57(±0.02) 0.54(±0.06) QED(↓) Single-modal 0.20(±0.04) 0.15(±0.05) 0.15(±0.06) Multi-modal 0.20(±0.08) 0.12(±0.03) 0.12(±0.02) weight(↓) Single-modal 112.10(±20.66) 67.83(±15.92) 89.75(±37.98) Multi-modal 105.72(±38.35) 68.01(±4.04) 95.62(±9.33)
Table 4: Molecule generation performance comparison on MOSES-MM.
Datasets Validity(%)↑ Uniqueness(%)↑ Novelty(%)↑
QM9 48 75.17 (±1.43) 64.13 (±1.58) 89.43 (±0.01)
Pcdes 29 84.31 (±0.24) 82.64 (±1.60) 93.253 (±0.59)
PubChemSTM 28 82.24 (±0.70) 80.89 (±1.12) 95.91 (±1.93)
igcdata 23 78.34 (±2.22) 61.21 (±1.50) 95.07 (±0.07)
Ours 84.8 (±0.6) 93.86 (±1.49) 96.51 (±1.57)
Table 5: Model performance comparison among different datasets.
Finally, we compare the molecule generation performance of GPT-3.5 across various datasets: QM9, Pcdes, PubChemSTM, igcdata, and our M3-20M, focusing on Validity, Uniqueness, and Novelty. Table 5 presents the experimental results.
Using our dataset, the model achieves the highest validity score of 84.8%, surpassing all other datasets, indicating the effectiveness of M3-20M in generating chemically accurate molecules. For uniqueness, our dataset again leads with 93.86%, significantly higher than Pcdes and PubChemSTM, demonstrating its capacity to support the generation of diverse molecular structures essential for discovering novel compounds. Our dataset also achieves the highest novelty score of 96.51%, followed by PubChemSTM at 95.91% and Pcdes at 93.25%, underscoring its effectiveness in generating unique molecular structures.
Overall, the comparison across Validity, Uniqueness, and Novelty highlights the superiority of our M3-20M dataset. It enables the generation of chemically valid molecules with high diversity and novelty, making it a valuable resource for model training and tuning in drug design and discovery.
5.2 Molecular Property Prediction
We refer to MoleculeNet 4 and divide the molecular property prediction task into regression task and classification task.
Regression tasks:
To test the effectiveness of our multi-modal dataset in support of molecule property prediction tasks, we conduct an ablation experiment on the QM9-MM dataset with three settings: using only the SMILES strings of molecules; using both the SMILES strings and 3D coordinates of molecules; and using the SMILES strings, 3D coordinates and the textual descriptions of molecules. We consider two regression tasks: predicting the dipole moment ($\mu$) and isotropic polarizability ($\alpha$) of the given molecules. We employ GPT-3.5, GPT-4, and GLM-4 APIs to infer the molecular properties respectively. For each molecule to be inferred, we provide the corresponding molecule information in the prompt based on the different versions of the settings and randomly select four additional molecule examples to help the LLMs better understand the chemical structure and the relation between the molecule information and the prediction ground truth.
The metric we use for the regression tasks is Mean Absolute Error (MAE), measuring the average magnitude of errors between predicted values and actual values. The smaller the MAE value, the more accurate the prediction. The formula of the MAE metric is as follows:
$\text{MAE}=\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|$ (1)
while $n$ is the number of samples, $y_{i}$ is predicted value, and $\hat{y}_{i}$ is the ground truth value.
The results in Table 6 show that providing more information with multi-modal data in the prompt can help LLMs better understand molecules and their chemical properties. Except for the task of predicting the dipole moment using GPT-4, all other models and tasks achieve the best results on the multi-modal dataset, which proves the effectiveness and high quality of our dataset.
SMILES SMILES+3D SMILES+3D+texts
$\mu$ $\alpha$ $\mu$ $\alpha$ $\mu$ $\alpha$
GPT-3.5 1.78 (±0.16) 39.50 (±2.16) 1.82 (±0.36) 35.37 (±1.53) 1.56 (±0.28) 34.74 (±1.60)
GLM-4 1.27 (±0.13) 12.82 (±0.77) 1.13 (±0.07) 14.53 (±0.67) 1.01 (±0.13) 12.05 (±0.48)
GPT-4 1.17 (±0.12) 13.71 (±0.08) 1.28 (±0.20) 28.9 (±0.84) 1.27 (±0.30) 9.50 (±1.65)
Table 6: Performance comparison of models using different modal data.
Classification tasks:
We evaluate the performance of single-modal and multi-modal datasets on various molecular property prediction tasks, focusing on accuracy mean (ACC Mean), accuracy variance (ACC Variance), and accuracy standard deviation (ACC Standard Deviation). The results are summarized in Table 7.
ACC Mean measures the average accuracy of the models. Multi-modal datasets show improvements over single-modal datasets. For example, the BACE-MM dataset’s ACC Mean increases from 0.5680 to 0.5780, and the BBBP-MM dataset sees an increase from 0.2280 to 0.2720. The ClinTox-MM dataset improves from 0.9260 to 0.9280, while the HIV-MM dataset shows a higher ACC Mean for single-modal data (0.9740) compared to multi-modal (0.9680). These results suggest that multi-modal datasets typically enhance model accuracy. ACC Variance indicates the variability in accuracy. The ClinTox-MM dataset’s variance decreases from 0.0005 to 0.0002 with multi-modal data, indicating more stable performance. ACC Standard Deviation provides insight into the distribution of accuracy values. The ClinTox-MM dataset sees a reduction from 0.0230 to 0.0130, suggesting enhanced consistency.
ACC Mean ACC Variance ACC Standard Deviation
Single-modal Multi-modal Single-modal Multi-modal Single-modal Multi-modal
BACE-MM 0.5680 0.5780 0.0014 0.0018 0.0377 0.0427
BBBP-MM 0.2280 0.2720 0.0005 0.0008 0.0217 0.0277
ClinTox-MM 0.9260 0.9280 0.0005 0.0002 0.0230 0.0130
HIV-MM 0.9740 0.9680 0.0002 0.0002 0.0134 0.0148
Table 7: Accuracy comparison between single-modal and multi-modal datasets for molecular property prediction tasks.
We compare the performance of single-modal and multi-modal datasets on various molecular property prediction tasks with the Tox21-MM dataset as the benchmark. The results are summarized in Table 8.
For ACC Mean, multi-modal datasets generally show an improvement. For example, in the NR-AR task, the ACC Mean increases from 0.9620 to 0.9680 with multi-modal data. Similarly, the NR-AhR task shows a rise from 0.8740 to 0.9020, and the NR-Aromatase task improves from 0.9620 to 0.9720. These results suggest that multi-modal datasets generally enhance accuracy. For ACC Variance, both the NR-AR and NR-PPAR-gamma tasks show equal variances of 0.0003 and 0.0002, indicating consistent performance. The NR-ER task significantly reduces variance from 0.0014 to 0.0003 with multi-modal data, highlighting improved stability. For ACC Standard Deviation, the SR-HSE task decreases significantly from 0.0295 to 0.0114 with multi-modal data, suggesting enhanced consistency.
ACC Mean ACC Variance ACC Standard Deviation
Single-modal Multi-modal Single-modal Multi-modal Single-modal Multi-modal
NR-AR 0.9620 0.9680 0.0003 0.0003 0.0164 0.0164
NR-AR-LBD 0.9800 0.9800 0.0002 0.0003 0.0141 0.0173
NR-AhR 0.8740 0.9020 0.0006 0.0009 0.0251 0.0303
NR-Aromatase 0.9620 0.9720 0.0001 0.0005 0.0110 0.0217
NR-ER 0.9140 0.9240 0.0014 0.0003 0.0378 0.0167
NR-ER-LBD 0.9580 0.9560 0.0003 0.0009 0.0179 0.0297
NR-PPAR-gamma 0.9780 0.9780 0.0002 0.0002 0.0148 0.0148
SR-ARE 0.9060 0.8680 0.0002 0.0005 0.0152 0.0217
SR-ATAD5 0.9680 0.9600 0.0005 0.0003 0.0228 0.0158
SR-HSE 0.9520 0.9560 0.0009 0.0001 0.0295 0.0114
SR-MMP 0.8820 0.8960 0.0008 0.0016 0.0277 0.0404
SR-p53 0.9500 0.9340 0.0005 0.0014 0.0235 0.0378
Table 8: Accuracy comparison between single-modal and multi-modal datasets on Tox21-MM.
Additionally, fine-tuning Llama3-8b with our data shows notable improvements. Results in Table 9 shows that fine-tuning Llama3-8b with our M3-20M dataset leads to notable improvements in model performance across various tasks. Specifically, the multi-modal results show significant gains over the single-modal baseline for the BACE, Tox21, HIV, and ClinTox datasets, with the most pronounced improvement seen on the BACE dataset, where performance increases from 0.844 to 0.887. This proves that our dataset has a wide range of application scenarios and practical value. In fact, our dataset can also be used for downstream tasks such as name prediction, reaction prediction, retrosynthesis, text-based molecule design, molecule captioning etc.
Table 9: Fine-tuning Llama3-8b with our M3-20M dataset
Dataset Single-modal Multi-modal (ours)
BBBP 0.959 0.959
BACE 0.844 0.887
Tox21 0.817 0.849
HIV 0.780 0.803
ClinTox 0.985 0.998
6 Conclusion
In this paper, we introduce M3-20M, a new and large-scale multi-modal molecular dataset containing over 20 million molecules. Extensive experiments demonstrate M3-20M’s effectiveness in boosting the performance of various models in molecule generation and property prediction. M3-20M’s integration of SMILES strings, 2D graphs, 3D structural data, physicochemical properties, and molecular textual descriptions provides a rich foundation for developing high-performance generative models for AI-driven drug design and discovery.
{acknowledgement}
This work was supported in part by the National Natural Science Foundation of China (NO. 62372326, No. 62172300).
7 Data and Software Availability Statement
The dataset is openly accessible for non-commercial use, hosted on a publicly available platform (https://github.com/bz99bz/M-3). This project also provides detailed dataset documentation and the workflow code discussed in the present paper.

### v2

@@
-M3-20M: A Large-Scale Multi-Modal Molecule Dataset for AI-driven Drug Design and Discovery
+
 Abstract
-This paper introduces M3-20M, a large-scale Multi-Modal Molecular dataset that contains over 20 million molecules.
-Designed to support AI-driven drug design and discovery, M3-20M is 71 times more in the number of molecules than the largest existing dataset, providing an unprecedented scale that can highly benefit training or fine-tuning large (language) models with superior performance for drug design and discovery.
-This dataset integrates one-dimensional SMILES, two-dimensional molecular graphs, three-dimensional molecular structures, physicochemical properties, and textual descriptions collected through web crawling and generated by using GPT-3.5, offering a comprehensive view of each molecule.
-To demonstrate the power of M3-20M in drug design and discovery, we conduct extensive experiments on two key tasks: molecule generation and molecular property prediction, using large language models including GLM4, GPT-3.5, and GPT-4.
+This paper introduces M3-20M, a large-scale Multi-Modal Molecule dataset that contains over 20 million molecules, with the data mainly being integrated from existing databases and partially generated by large language models.
+Designed to support AI-driven drug design and discovery, M3-20M is 71 times more in the number of molecules than the largest existing dataset, providing an unprecedented scale that can highly benefit the training or fine-tuning of models, including large language models for drug design and discovery tasks.
+This dataset integrates one-dimensional SMILES, two-dimensional molecular graphs, three-dimensional molecular structures, physicochemical properties, and textual descriptions collected through web crawling and generated using GPT-3.5, offering a comprehensive view of each molecule.
+To demonstrate the power of M3-20M in drug design and discovery, we conduct extensive experiments on two key tasks: molecule generation and molecular property prediction, using large language models including GLM4, GPT-3.5, GPT-4, and Llama3-8b.
 Our experimental results show that M3-20M can significantly boost model performance in both tasks.
-Specifically, it enables the models to generate more diverse and valid molecular structures and achieve higher property prediction accuracy than the existing single-modal datasets, which validates the value and potential of M3-20M in supporting AI-driven drug design and discovery.
+Specifically, it enables the models to generate more diverse and valid molecular structures and achieve higher property prediction accuracy than existing single-modal datasets, which validates the value and potential of M3-20M in supporting AI-driven drug design and discovery.
 The dataset is available at https://github.com/bz99bz/M-3.
 keywords:
-American Chemical Society,
-[…]
-{tocentry}
+Multi-modal Molecule Dataset, Dataset construction, Molecule generation, Molecule property prediction
 1 Introduction
-Drug design and discovery is an important field in pharmacy, which aims to identify new therapeutic compounds and optimize their properties for clinical use 1, 2, 3, 4, 5.
-This process involves the prediction of molecule-target interactions 6, 7, 8, 9 and molecular properties 10, 11, 12, 13, and the design or generation of targeted molecules 14, 15, 16, 17, 18, which are fundamental for developing effective and safe drugs.
-In recent years, various models trained or pre-trained and fine-tuned using massive data, such as 19, 20, 21, 22, 23, 24, 25, 26, 27, have emerged as powerful tools in this domain.
-These models leverage large datasets and sophisticated algorithms to generate new compounds, predict molecular structures, and analyze their biochemical properties, significantly advancing the process of drug design and discovery.
-However, most existing molecular datasets for model training are limited to single-modality, failing to capture the characteristics of molecules completely, thus hampering the building of powerful models for drug design and discovery.
-Recent works 23, 28, 29 have been paying increasing attention to multi-modal molecule datasets.
-Zeng 29 introduced a deep-learning system that bridges molecular structures and biomedical texts, achieving comprehension comparable to human professionals.
+As an important field in pharmacy, drug design and discovery aim to identify novel therapeutic compounds and optimize their properties for clinical application santos2017comprehensive ; hughes2011principles ; tsaioun2016evidence ; wu2018moleculenet ; chen2023artificial .
+This process involves a series of tasks, from the prediction of molecule-target interactions ji2023comprehensive ; ye2021unified ; li2022effective ; lu2024dynamicbind and molecular properties fu2024admetlab ; goldman2023prefix ; rong2020self ; masters2022gps++ , to the design or generation of targeted molecules gao2022sample ; franke2022probabilistic ; kong2022molecule ; yang2021hit ; song2024equivariant , which are fundamental for developing effective and safe drugs.
+In recent years, various models trained or pre-trained and fine-tuned using massive data, such as bagal2021molgpt ; chithrananda2020chemberta ; honda2019smiles ; fabian2020molecular ; liu2024git ; guo2023indeed ; liu2023group ; ai2024extracting ; yu2023multimodal ; ye2025drugassist , have emerged as powerful tools in this domain.
+These models leverage large datasets and sophisticated algorithms to generate new compounds, predict molecular structures, and analyze their biochemical properties, significantly accelerating drug design and discovery.
+However, most existing molecule datasets for model training are limited to single-modality, failing to capture molecular characteristics comprehensively, thus hindering the development of powerful models for drug design and discovery
+Recent works liu2024git ; liu2023multi ; zeng2022deep are paying increasing attention to multi-modal molecule datasets.
+Zeng zeng2022deep introduced a deep-learning system that bridges molecular structures and biomedical texts, achieving comprehension capacity comparable to human professionals.
 Liu et al.
-28 developed a multi-modal model for text-based retrieval and editing of molecular structures, facilitating advanced molecular text analysis.
+liu2023multi developed a multi-modal model for text-based retrieval and editing of molecular structures, facilitating advanced molecular text analysis.
 Liu et al.
-23 presented a multi-modal large language model for molecular science that integrates graphs, images, and texts to support various downstream tasks.
+liu2024git presented a multi-modal large language model for molecular science that integrates graphs, images, and texts to support various downstream tasks.
 However, datasets used in these works still suffer from significant limitations.
 On the one hand, these datasets are typically small in size, encompassing a limited chemical space that restricts the generalization power of the trained or tuned models.
-On the other hand, the absence of complete modalities of molecular data undermines the performance of computational models.
+On the other hand, the absence of complete modalities of molecular data undermines the performance of trained models.
 To address these limitations, this paper presents a new and large-scale multi-modal integrated molecule dataset called M3-20M for AI-driven drug design and discovery.
-M3-20M contains more than 20 million molecules with their SMILES strings, 2D graphs, 3D structures, physicochemical properties, and textual descriptions, providing unprecedented data scale, diversity, and comprehensiveness for training models of superior performance for various drug design and discovery downstream tasks.
-Concretely, M3-20M is 71 times more in the number of molecules than the largest existing dataset28.
-It collects physicochemical properties from the PubChem database 30 and enriches the description texts of molecules by both web crawling and GPT-3.5 generation.
+Data in M3-20M are mainly integrated from existing molecular databases, and supplemented by generation with large language models (LLMs).
+M3-20M contains more than 20 million molecules with their SMILES strings, 2D graphs, 3D structures, physicochemical properties, and textual descriptions, providing unprecedented data scale, diversity, and comprehensiveness for training models of drug design and discovery downstream tasks.
+M3-20M is 71 times more in the number of molecules than the largest existing multi-modal dataset PubChemSTM liu2023multi .
+It collects physicochemical properties from the PubChem database kim2016pubchem and enriches the description texts of molecules by web crawling and GPT-3.5 generation.
 Therefore, each molecule’s SMILES corresponds to its 2D graph and 3D structure, physicochemical properties, and description texts.
@@
 It also can aid in tasks such as pharmacokinetics modeling and drug-target interaction prediction.
-Moreover, we construct seven multimodal sub-datasets, QM9-MM, MOSES-MM, BACE-MM, BBBP-MM, HIV-MM, ClinTox-MM, and Tox21-MM, to facilitate more accurate and robust molecular property prediction.
-To demonstrate the power of M3-20M in supporting drug design and discovery, extensive experiments on two basic tasks, molecule generation and molecular property prediction, are conducted using large language models including GLM4 31, GPT-3.5 32, and GPT-4 33.
-The experimental results show that M3-20M can significantly boost model performance in both tasks.
-Concretely, with M3-20M, the above models can generate more diverse and valid molecular structures, and achieve higher property prediction accuracy than with existing single-modal datasets, thus verifying the value and potential of M3-20M in supporting AI-driven drug design and discovery.
+Moreover, we construct seven multi-modal sub-datasets, QM9-MM, MOSES-MM, BACE-MM, BBBP-MM, HIV-MM, ClinTox-MM, and Tox21-MM, to facilitate more accurate and robust molecular property prediction.
+To demonstrate M3-20M’s value and potential in supporting drug design and discovery, we conduct extensive experiments on two fundamental tasks, molecule generation and molecular property prediction.
+Our evaluation encompasses three categories of models, including closed-source LLMs (GLM4 du2022glm , GPT-3.5 ouyang2022training , and GPT-4 achiam2023gpt ), open-source LLM (Llama3-8b dubey2024llama ) and specialized molecule generation models (MoFlow zang2020moflow and D2L-OMP guo2024diffusing ).
+Our experimental results show that M3-20M can significantly boost model performance in both tasks.
+Concretely, with M3-20M, the above models can generate more diverse and valid molecular structures, and achieve higher property prediction accuracy than using existing single-modal datasets.
 Figure 1: Some examples from our M3-20M dataset.
 2 Related work
-Multi-modal molecular data is becoming increasingly important in artificial intelligence and molecular biology.
-which can provide more comprehensive and rich information, bringing new opportunities and challenges to drug discovery, molecular property prediction, and generation tasks.
-For example, Wang et al.
-34 pointed out that the use of multi-modal data, including sequence, structure, text etc., can improve the accuracy of drug discovery.
-However, the challenges faced by multi-modal data cannot be ignored, as different data types often necessitate distinct methods for organization, storage and processing as well as learning.
-Additionally, ensuring data consistency and integrity also poses a critical issue, with various forms of data potentially exhibiting semantic gaps or inconsistencies.
-Therefore, there is an urgent need to overcome the above issues to create a large-scale multi-modal integrated dataset for model training or tuning in drug design and discovery.
+Multi-modal molecular data offers comprehensive AI and molecular biology insights, enhancing drug discovery, property prediction, and generation tasks.
+As Wang et al.
+wang2023multitask demonstrated, integrating sequence, structure, and text data improves drug discovery accuracy.
+However, challenges persist in organization, processing, and maintaining consistency across different modalities.
+There’s an urgent need for large-scale integrated multi-modal datasets for drug design model development.
 In what follows, we briefly review the related work in three perspectives, molecular representation, single and two-modality molecule datasets, and multi-modality molecule datasets.
 2.1 Molecular representations
-Molecules can be represented by various modalities of information 35, 36, including one-dimensional sequences, two-dimensional graphs or images, three-dimensional structures, and molecular properties as well as textual descriptions.
-One-dimensional sequences primarily consist of SMILES strings, molecular fingerprints, and algebraic topology-based characterizations.
-Two-dimensional representations include molecular graphs and images generated using toolkits such as RDKit 37 or Open Babel 38.
-Three-dimensional structures are categorized into molecular graphs and grids 39, containing atomic coordinates that provide detailed spatial information.
-Molecular textual descriptions offer additional information derived from experiments.
+Molecules can be represented through diverse modalities david2020molecular ; ertl2010molecular : 1D sequences (SMILES, fingerprints, topological characterizations), 2D visualizations (graphs, RDKit/Open Babel landrum2013rdkit ; o2011open images), 3D structures (atomic coordinate graphs/grids li2022deep ), and textual descriptions from experiments.
 Single-modal data is a reduction of the complete structural information of the molecules, which means the representations learned from single-modal suffer from information loss.
@@
 2.2 Single-modal and two-modal molecule datasets
-Most existing molecule datasets are limited to single or dual modalities, restricting the generalization power of trained models 40, 41, 42, 43, 44.
-These datasets typically have only one-dimensional (1D) SMILES strings, or two-dimensional (2D) structures, or three-dimensional (3D) structures, and usually lack textual molecular descriptions.
-For instance, MOSES 45 comprises 4,591,276 molecules filtered by molecular weight, ranging from 250 to 350 Daltons, and contains solely one-dimensional SMILES representations.
-Similarly, ZINC 46, which serves as a repository of commercially available compounds for virtual screening, contains over 1.3 billion molecules obtained from 310 catalogs across 150 vendors, encompassing 2D graphs and, for the majority, 3D structures.
-Zeng et al.
-29 presented a deep learning system that connects molecular structures with biomedical texts, demonstrating a level of comprehension comparable to that of human experts.
+Existing molecule datasets are predominantly limited to single or dual modalities (1D SMILES strings, 2D structures, or 3D conformations), restricting the generalization power of trained models wang2023multimodal ; fink2007virtual ; wang2022deep ; axelrod2022geom ; yang2024quandb .
+For instance, MOSES polykovskiy2020molecular comprises 4,591,276 molecules with only SMILES representations, while ZINC irwin2005zinc includes over 1.3B molecules with 2D graphs and some 3D structures.
+Although recent advances like Zeng et al.
+zeng2022deep presented a molecular structure-text system.
+Beaini et al.
+beaini2023towards developed Graphium, a vast repository of 100M molecules with 13B labels, yet it focuses on multi-task learning rather than multi-modal representation.
 2.3 Multi-modal molecule datasets
-With the development of AI-driven drug design and discovery, multi-modal data have become increasingly important in recent years.
-Up to now, only a few datasets have been introduced to address this need, each with distinct purposes and contributions.
-Liu et al.
-28 constructed PubChemSTM, a dataset containing over 280,000 chemical structure–text pairs.
-Despite PubChemSTM’s ability to generalize novel biochemical concepts across various benchmarks, its dataset volume remains insufficient.
-Subsequently, the igcdata dataset 23 was proposed, which comprises 220,000 descriptions sourced from PubChem and ChEBI-20 47.
-Although these two datasets provide valuable resources, their relatively small scale restricts their utility for training high-performance models.
-Beyond scale, these two datasets also lack the diversity and completeness of molecular modalities necessary for robust model development.
-Issues such as sparse or overly brief textual descriptions further exacerbate this limitation, leading to suboptimal training outcomes of drug design and discovery models.
-To address the limitations of existing datasets and support the training or tuning of high-performance models for drug design and discovery, we present the M3-20M dataset, which significantly expands the scale and scope of multi-modal molecular data.
-Table 1 offers a qualitative comparison of our dataset against existing multi-modal datasets across seven dimensions, including the number of molecules contained, whether consisting of 3D structures, 2D graphs, PubChem CID, physicochemical properties and downstream task datasets, and the number of description texts.
-From Table 1, we can see that our dataset M3-20M has the largest number of molcuels and covers the most comprehensive molecular data modalities.
-Property QM948 PCdes29 PubChemSTM28 igcdata23 M3-20M (Ours)
+The growth of AI-driven drug design and discovery has elevated the importance of multi-modal molecular data.
+Current multi-modal datasets remain limited: PubChemSTM liu2023multi contains 280,000 structure-text pairs but lacks sufficient volume; igcdata liu2024git includes 220,000 descriptions from PubChem and ChEBI-20 gaulton2012chembl ; and ChEBI-20-MM liu2025quantitative provides 32,998 molecules as an evaluation benchmark rather than a training dataset.
+These datasets suffer from inadequate scale, limited modality coverage, and sparse textual descriptions.
+To overcome these limitations, we present the M3-20M dataset, which significantly expands the scale and scope of multi-modal molecular data.
+Table 1 compares our dataset with existing multi-modal datasets across seven dimensions: molecule count, 3D structures, 2D graphs, PubChem CID, physicochemical properties, downstream tasks, and description texts.
+M3-20M has the largest number of molecules and covers the most comprehensive molecular data modalities.
+Table 1: Comparison between our dataset M3-20M and existing multi-modal datasets.
+Property QM9ramakrishnan2014quantum PCdeszeng2022deep PubChemSTMliu2023multi igcdataliu2024git M3-20M (Ours)
 SMILES 134K 1.5K 280K 220K 20M
@@
 #Text descriptions $\times$ 1.5K 280K 220K 20M
-Table 1: Comparison between our dataset M3-20M and existing multi-modal datasets.
-3 Dataset Construction
-This section introduces the dataset construction process and some statistic information of the dataset.
-Molecular data is collected from existing databases PubChem, ZINC and QM9, and then processed to get multiple representations of molecules, including SMILES strings, 2D and 3D structures, physicochemical properties, and textual descriptions.
-Notably, we enhance molecular textual descriptions by generating using GPT-3.5 with an expert scoring mechanism to ensure that these descriptions are both scientifically accurate and practically useful.
-All the data are integrated into a comprehensive dataset with over 20 million molecules.
-Figure 2: The construction pipeline of our multi-modal molecular dataset M3-20M.
+3 Methodology
+This section outlines our dataset construction process and statistics.
+Molecular data is collected from existing databases PubChem, ZINC, and QM9ruddigkeit2012enumeration ; ramakrishnan2014quantum , processing them into multiple representations of molecules, including SMILES strings, 2D and 3D structures, physicochemical properties, and textual descriptions.
+Notably, we enhance molecular descriptions using GPT-3.5 with an expert scoring mechanism to ensure scientific accuracy.
+The final dataset integrates over 20 million molecules.
+Figure 2: The construction pipeline of our multi-modal molecule dataset M3-20M.
 3.1 Data Collection
-Figure 2 shows the pipeline for constructing our multi-modal molecular dataset M3-20M.
+Figure 2 illustrates our M3-20M construction pipeline.
 To obtain molecules and the corresponding molecular graphs and 3D structures, we first collect molecules from the PubChem, ZINC, and QM9 databases for their extensive molecular coverage and reliable data quality.
@@
 Textual descriptions serve as a bridge for large language models to comprehend multi-modal molecules.
-The remarkable recognition accuracy and robustness of the CLIP 49 model exemplify the benefits of effectively linking textual descriptions with other modality information of molecules.
-Similarly, since textual descriptions of molecules facilitate a deeper understanding of chemical information, we meticulously construct these descriptions in the dataset using three approaches: 1) For the molecules that have textual descriptions in PubChem, we directly extract the texts from PubChem.
+For textual descriptions, we employ three approaches: 1) For the molecules that have textual descriptions in PubChem, we directly extract the texts from PubChem.
 2) For the molecules that do not have textual descriptions but have physicochemical properties in PubChem, we transform these physicochemical property values into textual descriptions, which will be detailed later.
@@
 So we resort to text augmentation.
-To generate textual descriptions for molecules having physicochemical property values in PubChem, we extract 26 key physicochemical properties and their values from PubChem, including Molecular Weight, XLogP3, Hydrogen Bond Donor Count, Hydrogen Bond Acceptor Count, Rotatable Bond Count, Exact Mass, Monoisotopic Mass, Topological Polar Surface Area, Heavy Atom Count, Formal Charge, Complexity, Isotope Atom Count, Defined Atom Stereocenter Count, Undefined Atom Stereocenter Count, Defined Bond Stereocenter Count, Undefined Bond Stereocenter Count, Covalently-Bonded Unit Count, Canonicalization Status, Physical Description, Color/Form, Odor, Boiling Point, Melting Point, Flash Point, and Solubility Density.
+To generate textual descriptions for molecules having physicochemical property values in PubChem, we extract 26 key physicochemical properties and their values from PubChem (Table 10).
 Concretely, we search for these properties using the compound identifier (CID) in PubChem, ensuring all data complied with licensing requirements and utilizing the open interface provided by PubChem.
@@
 Here, “property name” indicates any of the 26 properties, and “specific value” means a concrete value of the property.
-Furthermore, we expand seven downstream task datasets with multi-modal information from our dataset by associating the PubChem CIDs of molecules, including molecule generation datasets MOSES and QM9, and molecular property prediction datasets such as BBBP, BACE, HIV, Tox21, and ClinTox.
-This expansion enables model enhancement with multi-modal information for both molecule generation and property prediction tasks.
-The expanded datasets are denoted as MOSES-MM, QM9-MM, BBBP-MM, BACE-MM, HIV-MM, Tox21-MM, and ClinTox-MM, respectively.
+Additionally, we expand seven downstream task datasets (MOSES, QM9, BBBP, BACE, HIV, Tox21, ClinTox) with multi-modal information and the PubChem CIDs, creating their ”-MM” counterparts to enhance both molecule generation and property prediction tasks.
 3.2 Text Description Generation by GPT-3.5
-As mentioned above, for molecules lacking physicochemical properties or not present in PubChem, we generate molecular textual descriptions as a molecule captioning task by using the GPT-3.5 API.
+For molecules lacking physicochemical properties or not present in PubChem, we generate textual descriptions using GPT-3.5.
 Figure 3 illustrates the process of molecular description text generation.
-Initially, we use the following prompt to guide GPT-3.5 in generating descriptions based solely on SMILES strings: “You are an expert chemist.
-Given the molecular SMILES, your task is to provide a detailed description of the molecule using your experienced chemical knowledge.”
+Initially, we used a basic prompt identifying GPT-3.5 as an expert chemist, but this produced errors in SMILES interpretation and functional group identification.
 Figure 3: Molecular text description generation by GPT-3.5.
-However, this approach results in issues such as incorrect naming of SMILES strings and misidentifying functional groups.
-To address these problems, inspired by the CO-STAR (Context-Objective-Style-Tone-Audience-Response) framework, we rewrite the original prompt as follows to enhance the output quality: “You are an expert chemist with a comprehensive understanding of molecular structures and chemistry.
+Inspired by the CO-STAR (Context-Objective-Style-Tone-Audience-Response) framework, we develop an enhanced prompt: “You are an expert chemist with a comprehensive understanding of molecular structures and chemistry.
 Using your extensive knowledge, please describe the molecule represented by the following SMILES string.
@@
 Use only the information inferred from the SMILES representation; do not create or infer any hypothetical or novel names for the molecule.”
-Utilizing this enhanced prompt, we generated 1,073,845 descriptions using GPT-3.5, which is only 0.934% of the total descriptions.
+Utilizing this enhanced prompt, we generated 1,073,845 descriptions, which is only 0.934% of the total descriptions.
 The proportion of synthetic data to real data is very low and will not have a negative impact on real data.
-The following section elaborates on the quality control measures implemented through the Human Expert Scoring Mechanism to ensure the precision and reliability of the generated descriptions.
+The subsequent section details our Human Expert Scoring Mechanism for quality control.
 3.3 Generated Text Quality Control by Human Expert Scoring
-To ensure the reliability and utility of the generated textual descriptions, we implement an expert scoring mechanism inspired by some recent works 50, 51.
-We recruit six undergraduate and postgraduate chemistry students from top China universities, all having excellent academic records.
-They evaluate the accuracy and usefulness of the generated texts, ensuring that the textual descriptions are both scientifically valid and practically useful.
-This mechanism assesses the textual descriptions from four dimensions: accuracy, effectiveness, comprehensiveness, and simplicity, as established in previous works on natural language generation and evaluation 52, 53.
+To ensure high-quality textual descriptions, we design a quality control protocol inspired by some recent works mccann2018natural ; kwiatkowski2019natural , which consists of three parts: selection and training of evaluation experts (or validators), scoring criteria, and stratified random sampling.
+3.3.1 Evaluation Expert Selection and Training
+To evaluate the quality of generated texts, six chemistry specialists (3 undergraduate and 3 postgraduate students) from top Chinese universities with documented expertise in molecular characterization were recruited as validators.
+All validators received a standardized training program covering evaluation criteria, error classification, and scoring calibration using 200 pilot samples.
+3.3.2 Quality Assessment Criteria
+This mechanism assesses the textual descriptions from four dimensions: accuracy, effectiveness, comprehensiveness, and simplicity, as established in previous works on natural language generation and evaluation wang2018glue ; post2018call .
 Each description can earn a maximum score of 10 points.
@@
 Descriptions with significant inaccuracies (0 points in accuracy) will also be regenerated.
-4 Statistic Results
-This section presents some statistic information of the M3-20M dataset to highlight its comprehensive annotations, diverse textual descriptions and physicochemical properties, which can significantly enhance its utility for molecular research and AI-driven drug discovery.
-By comparing it with existing datasets such as PubChem, we emphasize the enhancements in both scale and quality that M3-20M offers.
-The section further delves into word distribution — a key feature of textual data, and explores correlations among physicochemical properties through advanced statistical techniques like Principal Component Analysis (PCA).
-Additionally, we outline the robust maintenance protocols established to ensure our dataset’s accuracy and usability for the research community.
-These findings demonstrate the dataset’s value as a fundamental resource for molecular representation learning and related applications.
-4.1 Textual Descriptions
-Figure 4 shows a comparison of molecule descriptions between PubChem and our M3-20M.
-The PubChem description bar is relatively small, reaching only 360,133, while the M3-20M annotation description bar towers over it, marking a count of 20,249,090.
-This enhancement is critical for AI-driven drug design and discovery, as the increased volume and details of descriptions facilitate more comprehensive analysis and model training.
-Figure 4: Comparison of molecule descriptions: PubChem vs. our M3-20M
-4.2 Word Distribution in Description Texts
-Figure 5 presents the top 20 most frequently occurring words in the textual descriptions of our dataset, plotted alongside their respective frequencies to emphasize their relative prominence.
-The most commonly used terms, such as ’natural,’ ’product,’ and ’acid,’ appear tens of thousands of times, underscoring their central role in describing molecular properties and functions.
-Additional notable terms, including ’related,’ ’organisms,’ and ’amino,’ further highlight the biological and chemical contexts associated with these molecules.
-This visualization of word frequencies provides valuable insights into molecular datasets’ linguistic patterns and descriptive focus.
-Figure 5: Top 20 most frequent words in textual descriptions.
-4.3 Physicochemical Properties
-Table 2 and Figure 6 provide a detailed overview of the specific names, corresponding quantities, and reference of the 26 key physicochemical properties extracted from PubChem.
-Of the 26 key physicochemical properties, 18 are computed using professional software  [^1] PubChem 2.2; Cactvs 3.4.8.18, while the remaining 8 represent experimental values sourced from the CAMEO Chemicals database and Hazardous Substances Data Bank (HSDB).
-Notably, 18 of these properties exhibit quantities reaching tens of millions, underscoring our dataset’s extensive coverage and comprehensiveness.
-Physicochemical Properties Number Reference Molecular Weight 19174000 Computed by PubChem 2.2 XLogP3 16340203 Computed by XLogP3 3.0 Hydrogen Bond Donor Count 19173687 Computed by Cactvs 3.4.8.18 Hydrogen Bond Acceptor Count 19173559 Computed by Cactvs 3.4.8.18 Rotatable Bond Count 19173464 Computed by Cactvs 3.4.8.18 Exact Mass 19173398 Computed by PubChem 2.2 Monoisotopic Mass 19173327 Computed by PubChem 2.2 Topological Polar Surface Area 19173287 Computed by Cactvs 3.4.8.18 Heavy Atom Count 19173259 Computed by PubChem Formal Charge 19173222 Computed by PubChem Complexity 19173210 Computed by Cactvs 3.4.8.18 Isotope Atom Count 19173193 Computed by PubChem Defined Atom Stereocenter Count 19173170 Computed by PubChem Undefined Atom Stereocenter Count 19173153 Computed by PubChem Defined Bond Stereocenter Count 19173147 Computed by PubChem Undefined Bond Stereocenter Count 19173137 Computed by PubChem Covalently-Bonded Unit Count 19173132 Computed by PubChem Compound Is Canonicalized 19173130 Computed by PubChem Physical Description 2031 CAMEO Chemicals database Color/Form 246 Hazardous Substances Data Bank (HSDB) Odor 100 Hazardous Substances Data Bank (HSDB) Boiling Point 130 CAMEO Chemicals database Melting Point 633 CAMEO Chemicals database Flash Point 35 CAMEO Chemicals database Solubility 10813 CAMEO Chemicals database Density 170 CAMEO Chemicals database
-Table 2: The 26 key physicochemical properties and their corresponding numbers and references.
-Figure 6: Physicochemical property distribution.
-Figure 7: Correlation matrix of physicochemical properties.
-Figure 8: Visualization of molecular space with principal component analysis (PCA).
-As shown in Figure 7, we first calculate the correlation matrix of all physicochemical properties and draw a heat map.
-We find that Molecular Weight and XLogP3 are the most suitable because they exhibit strong correlations with several other key properties, indicating their ability to capture significant variance in the dataset.
-Thus, we select these two properties as the first and second principal components for Principal Components Analysis (PCA) and visualize the molecular space of randomly sampled 10000 molecules as shown in Figure 8.
-4.4 Dataset Maintenance
-To ensure the accessibility and accuracy of the M3-20M dataset, we have organized a dedicated maintenance team consisting of six members, including three doctoral students and three master’s students.
-This team will perform regular maintenance twice a week, correcting any errors and updating the dataset as needed.
-This rigorous maintenance protocol supports the community’s exploration of molecular representation and downstream molecular tasks and facilitates the training of large models with improved versatility and robustness.
-4.5 Supporting Various Downstream Tasks
-The M3-20M dataset supports model training and tuning for various downstream drug design and discovery tasks.
-While general large language models have achieved remarkable success in natural language processing, their performance in specialized fields like drug design and discovery still has significant room for improvement.
-Figure 9 illustrates how to use the M3-20M dataset to enhance large language models (LLMs) or train models for downstream drug design and discovery tasks, which consists of three paradigms: prompting, finetuning, and training from scratch.
-Prompt engineering involves carefully constructing input prompts to guide a general large language model (LLM) in generating more accurate outputs tailored to drug design and discovery.
-By leveraging our M3-20M dataset, prompt engineering allows the model to quickly adapt to this field without altering its internal parameters.
-Fine-tuning refers to further training the LLMs on the M3-20M dataset, adjusting certain model parameters based on existing pre-training.
-This process enables the model to better align with the specific characteristics and needs of drug design and discovery while retaining its general language knowledge.
-Training from scratch entails building and training a language model on drug design and discovery domain data.
-This approach necessitates a large volume of training data, which our M3-20M dataset can adequately provide.
-By utilizing our M3-20M dataset through prompt engineering, fine-tuning, and training from scratch, the general LLM’s drug design and discovery capabilities can be enhanced from multiple perspectives and to varying degrees.
-M3-20M also can aid in tasks such as pharmacokinetics modeling and drug-target interaction prediction.
-Moreover, we construct seven multimodal sub-datasets, QM9-MM, MOSES-MM, BACE-MM, BBBP-MM, HIV-MM, ClinTox-MM, and Tox21-MM, to facilitate more accurate and robust molecular property prediction.
-We also provide tools for generating molecular images to support compound patent searches and for crawling PubMed to find articles related to specific molecules, thereby enhancing textual descriptions.
+3.3.3 Sampling Strategy and Validation Results
+We adopt stratified random sampling to sample 12,000 molecules (2,000 per validator) from our dataset, ensuring no overlap between validators.
+The validation reveals that the qualification rate is 98.8% (11,856/12,000), i.e., with only 144 instances requiring regeneration (24 ± 3.1 per validator).
+All the regenerated text descriptions underwent a secondary validation.
+This rigorous validation ensures that $\geq$ 99% of released descriptions meet strict scientific standards.
+3.4 Multi-Modal Datasets for Downstream Tasks
+To support model training for downstream tasks, we construct multi-modal datasets by integrating PubChem CIDs and descriptive text, thereby providing richer semantic context and molecular identity.
+In particular, we derive seven additional datasets from M3-20M, they are MOSES-MM and QM9-MM for molecule generation, while QM9-MM, BBBP-MM, BACE-MM, HIV-MM, Tox21-MM, and ClinTox-MM for molecular property prediction.
+Table 2 summarizes these datasets.
+Table 2: Statistics of the multi-modal datasets this work augmented for downstream tasks, including the number of tasks, the number of molecules, task type, and evaluation metrics
+Dataset #Tasks #Molecules Task type Metric
+MOSES-MM 1 1,936,962 Generation SNN/Test, Frag/Test, Scaf/Test, Filters, logP, SA, QED, weight
+QM9-MM 2 6,830 Generation and Regression Validity, Uniqueness, Novelty, MAE
+BBBP-MM 1 2,039 Classification ROC–AUC
+BACE-MM 1 1,513 Classification ROC–AUC
+ClinTox-MM 2 1,478 Classification ROC–AUC
+Tox21-MM 12 7,831 Classification ROC–AUC
+HIV-MM 1 41,127 Classification ROC–AUC
+4 Experiments
+In this section, we evaluate the value and potential of the M3-20M dataset by experiments of two fundamental downstream tasks: molecule generation and molecular property prediction.
+In the experiments, we employ two different model adaptation approaches: (1) Few-shot prompting with in-context learning (FP-ICL) brown2020language on closed-source models (GLM4, GPT-3.5, and GPT-4), where we design input prompts with demonstration examples from M3-20M to guide the models’ responses without modifying model parameters; (2) Parameter-efficient fine-tuning using low-rank adaptation of large language models (LoRA) hu2021lora on the open-source Llama3-8b model, which updates a small subset of model parameters while maintaining the base model’s knowledge.
+Additionally, we present comprehensive statistical analyses on M3-20M to demonstrate its potential for supporting downstream drug design and discovery tasks.
+Our findings show that M3-20M can significantly improve model performance in the two molecular tasks with both adaptation approaches.
+4.1 LLM Adaptation Methods
+Large language models (LLMs) can be adapted for domain-specific tasks through various methods.
+As illustrated in Fig.
+4, we employ two methods that differ significantly in their resource requirements and implementation.
+4.1.1 Few-shot Prompting with In-context Learning (FP-ICL)
+This approach involves providing carefully selected example pairs from M3-20M alongside the target task in a single prompt.
+By incorporating these demonstrations, the model learns to recognize patterns and generate appropriate responses for new cases without any parameter update.
+This method requires only API access to LLMs (such as GPT-4, GPT-3.5, and GLM4) and minimal computational resources, making it highly accessible and cost-effective for practical applications.
+4.1.2 Parameter-efficient Fine-tuning with LoRA
+Low-Rank Adaptation (LoRA) represents a more sophisticated adaptation method that modifies a subset of model parameters while preserving the model’s base knowledge.
+This approach requires access to open-source models (such as Llama3-8b) and substantial computational resources (e.g.
+RTX4090 GPU).
+While more resource-intensive, LoRA enables deeper model adaptation for specialized tasks.
+While M3-20M is also suitable for training LLMs from scratch, such training requires powerful computational infrastructure beyond the scope of this study.
+Our adoption of these two adaptation methods provides practical insights into the dataset’s utility under different resource constraints.
+4.2 Molecule Generation Experiments
+The M3-20M dataset enables de novo molecule generation by providing rich contextual information that improves the understanding of chemical structures and properties of generative models.
+4.2.1 Experimental Details
+We employ the FP-ICL method to leverage multi-modal information for molecular generation by LLMs.
+Both general-purpose closed-source LLMs (GLM-4, GPT-3.5, GPT-4) and specialized generative models (MoFlow and D2L-OMP) are used.
+We compare the performance of using multi-modal molecules and single-modal molecules as examples for generation.
+For the multi-modal examples, we provide 10 randomly selected molecules from our M3-20M dataset with their SMILES strings, 3D coordinates, and textual descriptions to the LLMs.
+The single-modal examples use only SMILES strings.
+100 molecules per configuration are generated for performance assessment.
+We trained these models using QM9-MM to maintain compatibility with their existing network structures and parameters.
+Both models are trained on RTX4090 with 24GB memory and batch size is 32.
+The flow-based model (MoFlow) is trained for 200 epochs, while the diffusion-based model (D2L-OMP) is trained for 300 epochs.
+The rest of the parameters of the two models are the same as those in the original paper.
+We perform each experiment five times to ensure statistical reliability and report the averaged results and variance.
+We employ the FP-ICL method to leverage multi-modal information for molecular generation by LLMs.
+Both general-purpose closed-source LLMs (GLM-4, GPT-3.5, GPT-4) and specialized generative models (MoFlow and D2L-OMP) are used.
+We compare the performance of using multi-modal molecules and single-modal molecules as examples for generation.
+For the multi-modal examples, we provide 10 randomly selected molecules from our M3-20M dataset with their SMILES strings, 3D coordinates, and textual descriptions to the LLMs.
+The single-modal examples use only SMILES strings.
+100 molecules per configuration are generated for performance assessment.
+We train MoFlow and D2L-OMP using QM9-MM on RTX4090 with 24GB memory with the batch size being 32.
+MoFlow is trained with 200 epochs, while D2L-OMP is trained with 300 epochs.
+The other parameters of these two models are the same as those in the original paper.
+We perform each experiment five times to ensure statistical reliability and report the averaged results and variance.
 Figure 4: Paradigms of model training and finetuning using our dataset.
-5 Dataset Evaluation
-In this section, we evaluate the effectiveness of the M3-20M dataset on two fundamental downstream tasks: molecule generation and molecular property prediction.
-We primarily perform prompt tuning and in-context learning on the closed-source large language models GLM4, GPT-3.5, and GPT-4, while also conducting fine-tuning on the open-source model Llama3-8b.
-Additionally, our dataset can be applied to other tasks, including name prediction, reaction prediction, retrosynthesis, text-based molecule design, and molecule captioning.
-Our findings demonstrate that M3-20M significantly enhances model performance across these molecular tasks.
-5.1 Molecule Generation
-We utilize the GLM-4, GPT-3.5, and GPT-4 APIs to evaluate the power of M3-20M in molecule generation.
-Through in-context learning (ICL) 54, we provide 10 randomly selected multi-modal examples (including SMILES strings of molecules, three-dimensional atomic coordinates, and descriptions) and single-modal examples (containing only SMILES), respectively.
-Subsequently, 100 molecules are generated randomly for each setup.
+4.2.2 Performance Metrics
 We analyze the performance of models trained on single-modal and multi-modal datasets in molecule generation using three metrics: validity, uniqueness, and novelty.
-The results, presented in Table 3, demonstrate the superiority of multi-modal datasets across all metrics and models.
-Validity measures the proportion of chemically valid generated molecules.
-Multi-modal datasets consistently outperform single-modal ones.
-For instance, GLM4’s validity improves from 72.73% to 85.37%, GPT-3.5’s from 76.01% to 84.8%, and GPT-4’s from 92.3% to 97.99%.
-Uniqueness assesses the diversity of generated molecules.
-Multi-modal datasets show significant improvements: GLM4’s uniqueness rises from 68.17% to 82.25%, GPT-3.5’s from 46.46% to 93.86%, and GPT-4’s from 58.64% to 70.17%.
-Novelty evaluates the proportion of generated novel molecules.
-Multi-modal datasets show substantial improvements, with GLM4’s novelty increasing from 97.90% to 98.10%, GPT-3.5’s from 85.72% to 96.51%, and GPT-4’s from 90.66% to 98.95%.
+Validity measures the proportion of chemically correct molecular structures generated.
+Uniqueness reflects the diversity of generated molecules.
+Novelty is defined as the fraction of valid molecules that are not present in the training dataset, revealing the model’s ability to propose new molecular structures.
+Furthermore, we validate our results using the comprehensive MOSES benchmark with additional structural metrics.
+Similarity to the nearest neighbor (SNN/Test) measures reference manifold matching precision.
+Fragment similarity (Frag/Test) assesses BRICS fragment degen2008art distribution alignment in generated and reference sets.
+Scaffold similarity (Scaff/Test) evaluates Bemis-Murcko scaffold bemis1996properties distribution correspondence in generated and reference sets.
+Filters is the fraction of generated molecules that pass custom medicinal chemistry filters (MCFs) and PAINS filters baell2010new .
+The MOSES benchmark includes key chemical property metrics (logP, SA, QED, Molecular weight (MW)).
+We calculate Wasserstein-1 distances between generated and test set distributions for these properties.
+Lower distances indicate better alignment with realistic molecular characteristics, further confirming generation quality.
+4.2.3 Experimental Results
+The results in Table 3 demonstrate that models leveraging multi-modal datasets consistently outperform their single-modal counterparts using single-modal datasets across all metrics and model types.
+For validity, both specialized molecule generation models and general LLMs show remarkable performance improvement with multi-modal data.
+For uniqueness, the consistently high uniqueness scores across all models demonstrate multi-modal datasets boost their robust capability to generate diverse molecular structures.
+For novelty, our multi-modal dataset significantly enhances the performance of both LLMs and specialized molecule generation models.
+Table 3: Molecule generation performance comparison: multi-modal data vs. single-modal data.
 Validity(%)↑ Uniqueness(%)↑ Novelty(%)↑
 Single-modal Multi-modal Single-modal Multi-modal Single-modal Multi-modal
+Specialized Molecule Generation Models
+MoFlow 95.36($\pm$1.2) 96.21($\pm$1.24) 98.65($\pm$0.6) 98.71($\pm$0.19) 94.72($\pm$0.8) 99.12($\pm$0.13)
+D2L-OMP 98.60 ($\pm$0.83) 98.88($\pm$0.92) 99.80($\pm$0.17) 99.80($\pm$0.14) 83.39($\pm$0.52) 89.10($\pm$0.51)
+General LLMs
 GLM4 72.73($\pm$1.73) 85.37($\pm$1.58) 68.17($\pm$2.94) 82.25($\pm$2.95) 97.90($\pm$0.98) 98.10($\pm$1.23)
@@
 GPT-4 92.3($\pm$3.85) 97.99($\pm$0.41) 58.64($\pm$0.36) 70.17($\pm$4.92) 90.66($\pm$1.14) 98.95($\pm$0.34)
-Table 3: Molecule generation performance comparison: multi-modal data vs. single-modal data.
-Furthermore, Table 4 summarizes the molecular generation performance of our M3-20M dataset using MOSES-MM as the benchmark.
-Multi-modal data enhances SNN/Test performance for GPT-3.5 and GPT-4, showing increased similarity to the test set.
-For Frag/Test, multi-modal data boosts GLM4 to 0.27% and GPT-4 to 0.20%.
-In Scaf/Test, all models improve with multi-modal data; GLM4 rises to 0.07%, while GPT-3.5 and GPT-4 each improve to 0.01%.
-Regarding Filters, GLM4’s pass rate increases to 0.80% with multi-modal data, while GPT-3.5 maintains 1.0%.
-Metrics like logP, SA, QED, and molecular weight are assessed using the Wasserstein-1 distance to the MOSES-MM test set, where lower values indicate better performance.
-Multi-modal data significantly reduces logP distance, especially for GPT-3.5 (0.99) and GPT-4 (0.92), indicating superior predictions.
-In SA, multi-modal data improves scores for GPT-3.5 and GPT-4, reducing distances to 0.57 and 0.54, respectively.
-QED scores improve with multi-modal data, reaching 0.12 for both models.
-For molecular weight, predictions improve for GLM4, reducing the distance to 105.72.
-These findings demonstrate that our multi-modal dataset significantly enhances molecule generation models across nearly all metrics, highlighting the benefit of integrating diverse molecular representations.
-The results support the widespread use of multi-modal data in molecular research to develop more robust and versatile generative models.
-GLM4 GPT-3.5 GPT-4 SNN/Test(%)↑ Single-modal 0.49(±0.03) 0.53(±0.09) 0.49(±0.07) Multi-modal 0.44(±0.09) 0.56(±0.04) 0.52(±0.01) Frag/Test(%)↑ Single-modal 0.24(±0.02) 0.44(±0.12) 0.19(±0.08) Multi-modal 0.27(±0.14) 0.32(±0.01) 0.20(±0.01) Scaf/Test(%)↑ Single-modal 0.04(±0.05) 6.90e-4(±6.90e-4) 2.01e-3(±4.13e-4) Multi-modal 0.07(±0.09) 0.01(±0.01) 0.01(±4.99e-3) Filters(%)↑ Single-modal 0.73(±0.09) 0.8(±0.0) 1.0(±0.0) Multi-modal 0.80(±1.11e-16) 1.0(±0.0) 0.9(±0.10) logP(↓) Single-modal 1.35(±0.17) 1.63(±0.32) 1.27(±0.31) Multi-modal 1.44(±0.28) 0.99(±0.05) 0.92(±0.11) SA(↓) Single-modal 0.83(±0.22) 0.63(±0.11) 0.60(±0.14) Multi-modal 1.00(±0.10) 0.57(±0.02) 0.54(±0.06) QED(↓) Single-modal 0.20(±0.04) 0.15(±0.05) 0.15(±0.06) Multi-modal 0.20(±0.08) 0.12(±0.03) 0.12(±0.02) weight(↓) Single-modal 112.10(±20.66) 67.83(±15.92) 89.75(±37.98) Multi-modal 105.72(±38.35) 68.01(±4.04) 95.62(±9.33)
-Table 4: Molecule generation performance comparison on MOSES-MM.
+Furthermore, Table 4 summarizes the results of additional molecule generation performance metrics.
+Multi-modal training consistently enhances performance across nearly all indicators, enabling the model to exceed single-modal baselines in generating chemically meaningful structures.
+Improvements are evident in SNN, Frag, Scaff, and Filter scores.
+Multi-modal training also reduces Wasserstein-1 distance for logP, SA, and QED properties.
+These results demonstrate that the chemical properties of generated molecules align with those of the reference molecules, highlighting the effectiveness of multi-modal data in producing compounds with pharmacologically relevant characteristics while maintaining structural diversity and synthetic accessibility.
+Table 4: Molecule generation comparison using MOSES benchmark metrics.
+The table presents results of structural validity (SNN/Test, Frag/Test, Scaff/Test, Filters) and chemical property alignment (Wasserstein-1 distances for logP, SA, QED, MW) between using single-modal data and using multi-modal data.
+Lower Wasserstein-1 distance indicates better distribution matching with the reference compounds.
+GLM4 GPT-3.5 GPT-4 SNN/Test(↑) Single-modal 0.49(±0.03) 0.53(±0.09) 0.49(±0.07) Multi-modal 0.49(±0.09) 0.56(±0.04) 0.52(±0.01) Frag/Test(↑) Single-modal 0.24(±0.02) 0.32(±0.12) 0.19(±0.08) Multi-modal 0.27(±0.14) 0.44(±0.01) 0.20(±0.01) Scaf/Test(↑) Single-modal 0.04(±0.05) 6.90e-4(±6.90e-4) 2.01e-3(±4.13e-4) Multi-modal 0.07(±0.09) 0.01(±0.01) 0.01(±4.99e-3) Filters(↑) Single-modal 0.73(±0.09) 0.8(±0.0) 1.0(±0.0) Multi-modal 0.80(±1.11e-16) 1.0(±0.0) 1.0(±0.10) logP(↓) Single-modal 1.35(±0.17) 1.63(±0.32) 1.27(±0.31) Multi-modal 1.44(±0.28) 0.99(±0.05) 0.92(±0.11) SA(↓) Single-modal 0.83(±0.22) 0.63(±0.11) 0.60(±0.14) Multi-modal 1.00(±0.10) 0.57(±0.02) 0.54(±0.06) QED(↓) Single-modal 0.20(±0.04) 0.15(±0.05) 0.15(±0.06) Multi-modal 0.20(±0.08) 0.12(±0.03) 0.12(±0.02) Molecular weight(↓) Single-modal 112.10(±20.66) 67.83(±15.92) 89.75(±37.98) Multi-modal 105.72(±38.35) 68.01(±4.04) 95.62(±9.33)
+Table 5: Model performance comparison among different datasets.
 Datasets Validity(%)↑ Uniqueness(%)↑ Novelty(%)↑
-QM9 48 75.17 (±1.43) 64.13 (±1.58) 89.43 (±0.01)
-Pcdes 29 84.31 (±0.24) 82.64 (±1.60) 93.253 (±0.59)
-PubChemSTM 28 82.24 (±0.70) 80.89 (±1.12) 95.91 (±1.93)
-igcdata 23 78.34 (±2.22) 61.21 (±1.50) 95.07 (±0.07)
+QM9 ramakrishnan2014quantum 75.17 (±1.43) 64.13 (±1.58) 89.43 (±0.01)
+Pcdes zeng2022deep 84.31 (±0.24) 82.64 (±1.60) 93.253 (±0.59)
+PubChemSTM liu2023multi 82.24 (±0.70) 80.89 (±1.12) 95.91 (±1.93)
+igcdata liu2024git 78.34 (±2.22) 61.21 (±1.50) 95.07 (±0.07)
 Ours 84.8 (±0.6) 93.86 (±1.49) 96.51 (±1.57)
-Table 5: Model performance comparison among different datasets.
-Finally, we compare the molecule generation performance of GPT-3.5 across various datasets: QM9, Pcdes, PubChemSTM, igcdata, and our M3-20M, focusing on Validity, Uniqueness, and Novelty.
-Table 5 presents the experimental results.
-Using our dataset, the model achieves the highest validity score of 84.8%, surpassing all other datasets, indicating the effectiveness of M3-20M in generating chemically accurate molecules.
-For uniqueness, our dataset again leads with 93.86%, significantly higher than Pcdes and PubChemSTM, demonstrating its capacity to support the generation of diverse molecular structures essential for discovering novel compounds.
-Our dataset also achieves the highest novelty score of 96.51%, followed by PubChemSTM at 95.91% and Pcdes at 93.25%, underscoring its effectiveness in generating unique molecular structures.
-Overall, the comparison across Validity, Uniqueness, and Novelty highlights the superiority of our M3-20M dataset.
-It enables the generation of chemically valid molecules with high diversity and novelty, making it a valuable resource for model training and tuning in drug design and discovery.
-5.2 Molecular Property Prediction
-We refer to MoleculeNet 4 and divide the molecular property prediction task into regression task and classification task.
-Regression tasks:
-To test the effectiveness of our multi-modal dataset in support of molecule property prediction tasks, we conduct an ablation experiment on the QM9-MM dataset with three settings: using only the SMILES strings of molecules; using both the SMILES strings and 3D coordinates of molecules; and using the SMILES strings, 3D coordinates and the textual descriptions of molecules.
-We consider two regression tasks: predicting the dipole moment ($\mu$) and isotropic polarizability ($\alpha$) of the given molecules.
-We employ GPT-3.5, GPT-4, and GLM-4 APIs to infer the molecular properties respectively.
-For each molecule to be inferred, we provide the corresponding molecule information in the prompt based on the different versions of the settings and randomly select four additional molecule examples to help the LLMs better understand the chemical structure and the relation between the molecule information and the prediction ground truth.
-The metric we use for the regression tasks is Mean Absolute Error (MAE), measuring the average magnitude of errors between predicted values and actual values.
+Finally, we compare the molecule generation performance of GPT-3.5 using various datasets—QM9, Pcdes, PubChemSTM, igcdata, and our M3-20M—focusing on three key metrics: Validity, Uniqueness, and Novelty (see Table 5).
+The results demonstrate the advantages of our M3-20M dataset.
+For validity, GPT-3.5 consistently generates the most chemically accurate molecules using our dataset, underscoring M3-20M’s effectiveness in capturing essential chemical properties.
+Regarding uniqueness, the molecules produced with our data exhibit a greater diversity of structures, indicating the potential for discovering novel compounds.
+This trend continues with the novelty metric, that is, using our dataset can generate molecular structures distinct from those seen during training, further validating its ability to foster innovation in molecular design.
+Overall, the results above indicate that our dataset can significantly benefit molecule generation, offering a valuable resource for AI-driven molecule design and discovery.
+4.3 Molecular Property Prediction Experiments
+Accurate prediction of molecular properties is fundamental to drug discovery.
+To evaluate M3-20M’s effectiveness in property prediction tasks, we conduct extensive experiments following the settings in MoleculeNet wu2018moleculenet .
+Our evaluation encompasses regression and classification tasks, addressing diverse molecular properties crucial for drug development.
+We employ FP-ICL for regression tasks, both FP-ICL and parameter-efficient fine-tuning with LoRA for classification tasks.
+To prevent data leakage, we firstly partition datasets into non-overlapping training and test sets.
+Secondly, we mask all the target property-related information in textual descriptions during training and evaluation.
+We perform each experiment five times and report the averaged result and variance.
+4.3.1 Regression Tasks
+Experimental details.
+We apply FP-ICL for regression tasks on the QM9-MM dataset under three settings: using SMILES strings only, using both SMILES strings and 3D coordinates, and using SMILES strings with 3D coordinates plus textual descriptions.
+We focus on predicting dipole moment ($\mu$) and isotropic polarizability ($\alpha$).
+We use GPT-3.5, GPT-4, and GLM-4 APIs as LLMs, and randomly select four molecules from QM9-MM for the LLMs in the prompt according to the three specific settings.
+Performance metrics.
+The metric we use for the regression tasks is Mean Absolute Error (MAE), measuring the average error between predicted values and actual values.
 The smaller the MAE value, the more accurate the prediction.
-The formula of the MAE metric is as follows:
+Formally, MAE is evaluated as follows:
 $\text{MAE}=\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|$ (1)
-while $n$ is the number of samples, $y_{i}$ is predicted value, and $\hat{y}_{i}$ is the ground truth value.
+where $n$ is the number of samples, $y_{i}$ is predicted value, and $\hat{y}_{i}$ is the ground truth value.
+Experimental results.
 The results in Table 6 show that providing more information with multi-modal data in the prompt can help LLMs better understand molecules and their chemical properties.
 Except for the task of predicting the dipole moment using GPT-4, all other models and tasks achieve the best results on the multi-modal dataset, which proves the effectiveness and high quality of our dataset.
+Table 6: Comparison of molecular property prediction performance using multi-modal and single-modal data (regression task).
 SMILES SMILES+3D SMILES+3D+texts
@@
 GPT-4 1.17 (±0.12) 13.71 (±0.08) 1.28 (±0.20) 28.9 (±0.84) 1.27 (±0.30) 9.50 (±1.65)
-Table 6: Performance comparison of models using different modal data.
-Classification tasks:
-We evaluate the performance of single-modal and multi-modal datasets on various molecular property prediction tasks, focusing on accuracy mean (ACC Mean), accuracy variance (ACC Variance), and accuracy standard deviation (ACC Standard Deviation).
-The results are summarized in Table 7.
-ACC Mean measures the average accuracy of the models.
-Multi-modal datasets show improvements over single-modal datasets.
-For example, the BACE-MM dataset’s ACC Mean increases from 0.5680 to 0.5780, and the BBBP-MM dataset sees an increase from 0.2280 to 0.2720.
-The ClinTox-MM dataset improves from 0.9260 to 0.9280, while the HIV-MM dataset shows a higher ACC Mean for single-modal data (0.9740) compared to multi-modal (0.9680).
-These results suggest that multi-modal datasets typically enhance model accuracy.
-ACC Variance indicates the variability in accuracy.
-The ClinTox-MM dataset’s variance decreases from 0.0005 to 0.0002 with multi-modal data, indicating more stable performance.
-ACC Standard Deviation provides insight into the distribution of accuracy values.
-The ClinTox-MM dataset sees a reduction from 0.0230 to 0.0130, suggesting enhanced consistency.
+4.3.2 Classification Tasks
+Experimental details.
+First, we apply FP-ICL to regression tasks on BACE-MM, BBBP-MM, ClinTox-MM, HIV-MM, and Tox21-MM datasets with single-modal and multi-modal data respectively.
+Four randomly-selected molecules from each of the above datasets are used, with SMILES strings only for single-modal molecules, and SMILES strings, 3D coordinates and textual descriptions for multi-modal molecules.
+Then, we employ parameter-efficient fine-tuning with LoRA on Llama3-8b using our dataset M3-20M.
+Experiments are run on an RTX4090 with 24GB memory.
+LoRA hyperparameters include a rank of 4 and alpha of 16.
+Training lasts a maximum of 200 epochs with early stopping after 15-30 epochs without improvement.
+Performance metrics.
+We perform each experiment five times and report the averaged result and variance, using three metrics including accuracy mean (ACC Mean), accuracy variance (ACC Variance), and accuracy standard deviation (ACC Standard Deviation).
+These metrics provide comprehensive insight into prediction reliability and consistency across the different molecule datasets.
+Experimental results.
+Results from FP-ICL on classification tasks (Table 7) show using multi-modal data generally outperforms using single-modal data in molecular property prediction.
+On the majority of datasets including BACE-MM, BBBP-MM and ClinTox-MM, mean accuracy is improved, only on HIV-MM there is a slight performance decrease.
+Using multi-modal data usually yields reduced variance and standard deviation, particularly on the ClinTox-MM dataset, indicating enhanced prediction stability and consistency.
+These findings suggest that integrating multiple modalities provides complementary information and leads to more robust and accurate molecular property predictions.
+Table 7: Comparison of molecular property prediction performance using multi-modal and single-modal data (classification task).
 ACC Mean ACC Variance ACC Standard Deviation
@@
 HIV-MM 0.9740 0.9680 0.0002 0.0002 0.0134 0.0148
-Table 7: Accuracy comparison between single-modal and multi-modal datasets for molecular property prediction tasks.
-We compare the performance of single-modal and multi-modal datasets on various molecular property prediction tasks with the Tox21-MM dataset as the benchmark. The results are summarized in Table 8.
-For ACC Mean, multi-modal datasets generally show an improvement.
-For example, in the NR-AR task, the ACC Mean increases from 0.9620 to 0.9680 with multi-modal data.
-Similarly, the NR-AhR task shows a rise from 0.8740 to 0.9020, and the NR-Aromatase task improves from 0.9620 to 0.9720.
-These results suggest that multi-modal datasets generally enhance accuracy.
-For ACC Variance, both the NR-AR and NR-PPAR-gamma tasks show equal variances of 0.0003 and 0.0002, indicating consistent performance.
-The NR-ER task significantly reduces variance from 0.0014 to 0.0003 with multi-modal data, highlighting improved stability.
-For ACC Standard Deviation, the SR-HSE task decreases significantly from 0.0295 to 0.0114 with multi-modal data, suggesting enhanced consistency.
+Results on the Tox21-MM dataset are summarized in Table 8, from which we can still see that using multi-modal data can generally enhance molecular property prediction performance across multiple toxicity subtasks.
+On 8 of 12 subtasks we witness improved or equivalent accuracy, with notable improvements in NR-AhR and NR-Aromatase tasks.
+Using multi-modal data also frequently produces more consistent predictions, evidenced by the reduced variance and standard deviation in several subtasks, particularly NR-ER and SR-HSE.
+These results confirm that complementary information from multiple modalities enhances accuracy and reliability of toxicity prediction.
+Table 8: Comparison of molecular property prediction performance using multi-modal and single-modal data on Tox21-MM (classification task).
 ACC Mean ACC Variance ACC Standard Deviation
@@
 SR-p53 0.9500 0.9340 0.0005 0.0014 0.0235 0.0378
-Table 8: Accuracy comparison between single-modal and multi-modal datasets on Tox21-MM.
-Additionally, fine-tuning Llama3-8b with our data shows notable improvements.
-Results in Table 9 shows that fine-tuning Llama3-8b with our M3-20M dataset leads to notable improvements in model performance across various tasks.
-Specifically, the multi-modal results show significant gains over the single-modal baseline for the BACE, Tox21, HIV, and ClinTox datasets, with the most pronounced improvement seen on the BACE dataset, where performance increases from 0.844 to 0.887.
-This proves that our dataset has a wide range of application scenarios and practical value.
-In fact, our dataset can also be used for downstream tasks such as name prediction, reaction prediction, retrosynthesis, text-based molecule design, molecule captioning etc.
-Table 9: Fine-tuning Llama3-8b with our M3-20M dataset
+Table 9 presents the results of fine-tuning Llama3-8b with our M3-20M dataset.
+We also see notable improvements on model performance across various tasks.
+Concretely, using multi-modal data achieves significant gain over using single-modal data on BACE-MM, Tox21-MM, HIV-MM, and ClinTox-MM datasets, with the most pronounced improvement on the BACE dataset, where performance increases from 0.844 to 0.887.
+This shows that our dataset’s value and potential.
+Our dataset can also be used for other downstream tasks such as name prediction, reaction prediction, retrosynthesis, text-based molecule design, and molecule captioning, etc.
+Table 9: Fine-tuning Llama3-8b with our M3-20M dataset demonstrates significant performance improvement in terns of ACC mean.
 Dataset Single-modal Multi-modal (ours)
-BBBP 0.959 0.959
-BACE 0.844 0.887
-Tox21 0.817 0.849
-HIV 0.780 0.803
-ClinTox 0.985 0.998
+BBBP-MM 0.959 0.959
+BACE-MM 0.844 0.887
+Tox21-MM 0.817 0.849
+HIV-MM 0.780 0.803
+ClinTox-MM 0.985 0.998
+4.4 Statistic Analyses and Results
+This section presents statistical information on the M3-20M dataset to highlight its comprehensive annotations, diverse textual descriptions, and physicochemical properties, which can significantly enhance its utility for molecular research and AI-driven drug discovery.
+4.4.1 Textual Descriptions
+The PubChem description bar is relatively small, reaching only 360,133, while the M3-20M annotation description bar towers over it, marking a count of 20,249,090.
+This enhancement is critical for AI-driven drug design and discovery, as the increased volume and details of descriptions facilitate more comprehensive analysis and model training.
+4.4.2 Word Distribution in Description Texts
+Figure 5 shows the 20 most frequently occurring words in textual descriptions of our dataset, including “natural”, “product”, “acid”, and “metabolite” etc.
+These terms appear tens of thousands of times, reflecting the dataset’s emphasis on biologically derived compounds.
+The frequency distribution of descriptors like “conjugate”, “streptomyces” and “aza macrocycle” reveals the dataset’s focus on microbial secondary metabolites and their derivatives.
+This lexical analysis provides insight into the composition of the textual descriptions and highlights the predominance of natural product chemistry within the represented molecular space.
+Figure 5: 20 most-frequent words in textual descriptions.
+Table 10: The 26 key physicochemical properties and their corresponding numbers and references.
+Physicochemical Properties Number Reference
+Molecular Weight 19174000 Computed by PubChem 2.2
+XLogP3 16340203 Computed by XLogP3 3.0
+Hydrogen Bond Donor Count 19173687 Computed by Cactvs 3.4.8.18
+Hydrogen Bond Acceptor Count 19173559 Computed by Cactvs 3.4.8.18
+Rotatable Bond Count 19173464 Computed by Cactvs 3.4.8.18
+Exact Mass 19173398 Computed by PubChem 2.2
+Monoisotopic Mass 19173327 Computed by PubChem 2.2
+Topological Polar Surface Area 19173287 Computed by Cactvs 3.4.8.18
+Heavy Atom Count 19173259 Computed by PubChem
+Formal Charge 19173222 Computed by PubChem
+Complexity 19173210 Computed by Cactvs 3.4.8.18
+Isotope Atom Count 19173193 Computed by PubChem
+Defined Atom Stereocenter Count 19173170 Computed by PubChem
+Undefined Atom Stereocenter Count 19173153 Computed by PubChem
+Defined Bond Stereocenter Count 19173147 Computed by PubChem
+Undefined Bond Stereocenter Count 19173137 Computed by PubChem
+Covalently-Bonded Unit Count 19173132 Computed by PubChem
+Compound Is Canonicalized 19173130 Computed by PubChem
+Physical Description 2031 CAMEO Chemicals database
+Color/Form 246 Hazardous Substances Data Bank (HSDB)
+Odor 100 Hazardous Substances Data Bank (HSDB)
+Boiling Point 130 CAMEO Chemicals database
+Melting Point 633 CAMEO Chemicals database
+Flash Point 35 CAMEO Chemicals database
+Solubility 10813 CAMEO Chemicals database
+Density 170 CAMEO Chemicals database
+Figure 6: Visualization of molecular space with uniform manifold approximation and projection (UMAP).
+Figure 6 illustrates the diversity of our dataset in a two-dimensional space using UMAP mcinnes2018umap .
+UMAP (Uniform Manifold Approximation and Projection) excels as a dimensionality reduction technique by preserving both local and global structure while offering computational efficiency.
+It is particularly valuable for visualizing high-dimensional molecular fingerprints where topological relationships between compounds must be maintained.
+The UMAP projection of the M3-20M dataset reveals a predominantly interconnected chemical space with a dense central region and several peripheral clusters, suggesting substantial molecular diversity within a coherent chemical landscape.
+4.4.3 Dataset Maintenance
+A six-member maintenance team (three doctoral and three master’s students) will perform bi-weekly updates to the M3-20M dataset, correcting errors and implementing updates.
+This protocol enhances the dataset’s reliability for molecular representation research and supports more robust model training.
 5 Conclusion
-In this paper, we introduce M3-20M, a new and large-scale multi-modal molecular dataset containing over 20 million molecules.
+In this paper, we introduce M3-20M, a new and large-scale multi-modal molecule dataset containing over 20 million molecules.
 Extensive experiments demonstrate M3-20M’s effectiveness in boosting the performance of various models in molecule generation and property prediction.
-M3-20M’s integration of SMILES strings, 2D graphs, 3D structural data, physicochemical properties, and molecular textual descriptions provides a rich foundation for developing high-performance generative models for AI-driven drug design and discovery.
-{acknowledgement}
-This work was supported in part by the National Natural Science Foundation of China (NO. 62372326, No. 62172300).
+M3-20M’s integration of SMILES strings, 2D graphs, 3D structural data, physicochemical properties, and molecular textual descriptions provides a rich source for developing high-performance generative models for AI-driven drug design and discovery.
 6 Data and Software Availability Statement
-The dataset is openly accessible for non-commercial use, hosted on a publicly available platform (https://github.com/bz99bz/M-3).
-This project also provides detailed dataset documentation and the workflow code discussed in the present paper.
+The dataset is openly accessible under the GNU General Public License v3.0, hosted on a publicly available platform (https://github.com/bz99bz/M-3).
+7 Author contributions statement
+Siyuan Guo and Shuigeng Zhou conceived the experiment(s), Siyuan Guo, Lexuan Wang, Chang Jin, Jinxian Wang, Han Peng, and Huayang Shi conducted the experiment(s), Siyuan Guo, Lexuan Wang, and Chang Jin analysed the results.
+Siyuan Guo, Wengen Li, Jihong Guan, and Shuigeng Zhou wrote and reviewed the manuscript.
+8 Acknowledgments
+The authors thank the anonymous reviewers for their valuable suggestions. This work was supported in part by the National Natural Science Foundation of China (NO. 62372326, No. 62172300).
+Key Points M3-20M is the largest open-access multi-modal molecule dataset for AI-driven drug design and discovery, with over 20 million molecules.
+M3-20M has comprehensive modalities, integrating SMILES, 2D/3D structures, physicochemical properties, and textual descriptions.
+Experiments show that M3-20M can significantly boost the model performance in molecular generation and property prediction tasks.
+M3-20M supports diverse downstream tasks, including molecule generation, molecular property prediction, lead optimization, and virtual screening, etc.